### v1

# The ECME Algorithm Using Factor Analysis for DOA Estimation in Nonuniform Noise

Abstract
Maximum likelihood factor analysis has been used for direction of arrival estimation in unknown nonuniform noise and some iterative approaches have been developed. In particular, the Factor Analysis for Anisotropic Noise (FAAN) method proposed by Stoica and Babu has excellent convergence properties. In this letter, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the expectation-maximization algorithm, is designed, which has almost the same computational complexity at each iteration as the FAAN method. However, numerical results show that the ECME algorithm yields faster stable convergence and is computationally more efficient.
Index Terms:
Array signal processing, direction of arrival (DOA) estimation, expectation-maximization (EM) algorithm, factor analysis, maximum likelihood.
I Introduction
Factor analysis (FA) has been a tool in psychology for over one hundred years [1] and its purpose is to extract or estimate the low-rank covariance matrix from a measurement covariance matrix. To this end, several techniques have been developed, but the only one available, which provides efficient and consistent estimators, is maximum likelihood FA (MLFA) [2]. Unfortunately, these estimators cannot be expressed in closed form, which leads to a variety of iterative approaches [3]. In particular, one of the most promising is the Expectation-Maximization (EM) algorithm [4].
The EM algorithm regards observations or measurements as the incomplete data and designs the complete data containing missing (latent or unobserved) data. The complete data is so manipulable that its log-likelihood function (LLF) is more tractable than the incomplete-data or actual LLF. Hence, the EM algorithm uses the complete-data LLF to efficiently update the estimates of parameters, which can still increase the actual likelihood. To be specific, each iteration of the EM algorithm consists of an expectation step (E-step) and a maximization step (M-step). However, the M-step requires maximizing the complete-data LLF, which causes two shortcomings: 1) slow convergence, and 2) difficult maximization in numerous complex cases. To make up for both shortcomings, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the EM algorithm, has been proposed [5] and then applied to MLFA [6]. However, a conditional M-step (CM-step) at each iteration utilizes Newton-Raphson to update the estimates of partial parameters by the actual LLF and thus is complex as well as unstable.
In addition to psychometrics, econometrics, statistics and so on, FA recently finds applications in array signal processing, especially including direction of arrival (DOA) estimation in unknown nonuniform noise [7, 8]. The nonuniform noise assumption can let the signal received at an array be modeled by FA so that MLFA can be utilized to estimate the signal subspace for the subspace based DOA estimation [9]. For this purpose, an iterative approach, called Iterative ML Subspace Estimation (IMLSE), is proposed in [7]. This method sequentially updates the estimates of parameters at each iteration by explicit formulas and is computationally attractive. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [8]. Then, they propose a new iterative approach, called Factor Analysis for Anisotropic Noise (FAAN), with excellent convergence properties.
In this letter, the ECME algorithm is designed again for MLFA, which has almost the same computational complexity at each iteration as the FAAN method. However, numerical results show that the ECME algorithm yields faster stable convergence and is computationally more efficient. Importantly, the signal subspace estimated by the ECME algorithm can be used for the subspace based DOA estimation in unknown nonuniform noise.
Notations: $(\cdot)^{T}$ and $(\cdot)^{H}$ mean the transpose and conjugate transpose operators, respectively. $\mathrm{Det}(\boldsymbol{\Sigma})$, $\mathrm{Tr}\{\boldsymbol{\Sigma}\}$, $\boldsymbol{\Sigma}^{1/2}$, and $\boldsymbol{\Sigma}^{-1}$ stands for the determinant, trace, square root, and inverse of square matrix $\boldsymbol{\Sigma}$. $\Sigma_{i,j}$ is the element in the $i$th row and $j$th column of $\boldsymbol{\Sigma}$. $\boldsymbol{\Sigma}\succ\mathbf{0}_{N}$ and $\boldsymbol{\Sigma}\succeq\mathbf{0}_{N}$ denote that the $N\times N$ matrix $\boldsymbol{\Sigma}$ is positive definite and positive semi-definite, respectively. $\mathbf{I}_{N}$ is the identity matrix of order $N$ and $\jmath$ is the imaginary unit. $\mathcal{C}(\boldsymbol{\Sigma})$ denotes the column space of $\boldsymbol{\Sigma}$. $(\cdot)^{(k)}$ is an iterative value obtained at the $k$th iteration.
II Problem Description
We consider a uniform linear array (ULA) consisting of $N$ isotropic sensors for analytical simplicity. $M<N$ narrow-band source signals are assumed to impinge upon the array from far field and the distance between arbitrary two adjacent sensors is the half-wavelength of the source signals. Thus, the composite signal received at this ULA can be expressed as
$\mathbf{y}(t)=\sum_{m=1}^{M}\mathbf{a}(\theta_{m})s_{m}(t)+\mathbf{v}(t)=\mathbf{A}\mathbf{s}(t)+\mathbf{v}(t),$ (1)
where $\theta_{m}\in(0,\pi)$ represents the DOA of the $m$th source, $\mathbf{a}(\theta_{m})=[1~\cdots~e^{-(N-1)\jmath\pi\cos(\theta_{m})}]^{T}\in\mathbb{C}^{N}$ is the steering vector of the $m$th source, and $\mathbf{A}=[\mathbf{a}(\theta_{1})~\cdots~\mathbf{a}(\theta_{M})]$ is the array manifold matrix. Additionally, $\mathbf{s}(t)=[s_{1}(t)~\cdots~s_{M}(t)]^{T}\in\mathbb{C}^{M}$ is the source signal vector, $\mathbf{v}(t)$ means a spatially nonuniform Gaussian white noise vector with mean zero and covariance $\mathbf{Q}=\mathrm{diag}\{\sigma^{2}_{1},\dots,\sigma^{2}_{N}\}\succ\mathbf{0}_{N}$, i.e., $\mathbf{v}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{Q})$.
When the stochastic source signal model is adopted, $\mathbf{s}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{P})$, where $\mathbf{P}\succeq\mathbf{0}_{M}$ is the source covariance matrix. In (1), $\mathbf{s}(t)$ and $\mathbf{v}(t)$ are assumed to be mutually uncorrelated, which leads to $\mathbf{y}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{C})$ with $\mathbf{C}=\mathbf{A}\mathbf{P}\mathbf{A}^{H}+\mathbf{Q}$ [9]. In practice, we first sample $\mathbf{y}(t)$ and obtain independent and identically distributed snapshots as follows:
$\mathbf{y}(t)=\mathbf{A}\mathbf{s}(t)+\mathbf{v}(t),t=1,\dots,L$ (2)
where $L$ is the number of snapshots. Then, we obtain the DOA estimates $\hat{\theta}_{m}$’s by processing the snapshots $\mathbf{y}(t)$’s.
From the perspective of ML direction finding [10], we need to write the LLF of $\mathbf{Y}$ with $\mathbf{Y}=[\mathbf{y}(1)~\cdots~\mathbf{y}(L)]$ as
$l(\boldsymbol{\Phi})$ $=$ $\ln p(\mathbf{Y};\boldsymbol{\Phi})={\sum}_{t=1}^{L}\ln p\big{(}\mathbf{y}(t);\boldsymbol{\Phi}\big{)}$
$=$ $-LN\ln(\pi)-L\Big{(}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}\Big{)}$
where $\boldsymbol{\Phi}=(\boldsymbol{\theta},\mathbf{P},\mathbf{Q})$ with $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$, $p(\cdot)$ is the probability density function, and $\hat{\mathbf{R}}=(1/L)\sum_{t=1}^{L}\mathbf{y}(t)\mathbf{y}^{H}(t)\succ\mathbf{0}_{N}$ [^1] In this letter, we assume $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$, which requires $L\geq N$.. Importantly, we need to solve the equivalent ML direction finding problem [11, 12]:
$\min_{\boldsymbol{\theta}\in(0,\pi)^{M},\mathbf{P}\succeq\mathbf{0}_{M},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}.$ (3)
In order to solve problem (3), the ECME algorithm has been designed in the previous work [13]. In the next section, we will design the ECME algorithm using FA for the subspace based DOA estimation.
III MLFA Based Subspace Estimation
The ML direction finding problem (3) is intractable since $\mathbf{C}=\mathbf{APA}^{H}+\mathbf{Q}$ has excessive constraints on free parameters (unknowns) to be estimated. To alleviate this issue, we should release as many constraints as possible and introduce as many free parameters as possible. Thus, we aim not to solve problem (3) in this letter.
III-A MLFA Estimation Problem
It is well-known that the subspace based DOA estimation only requires the signal subspace $\mathcal{C}(\mathbf{A})$ [9]. As a result, let $\mathbf{C}$ be the population covariance matrix in FA [3]
$\mathbf{C}=\mathbf{S}\mathbf{S}^{H}+\mathbf{Q},$ (4)
where all elements in $\mathbf{S}\in\mathbb{C}^{N\times M}$ are free parameters (factor loadings) and hence $\mathbf{SS}^{H}$ must be able to equal any $\mathbf{APA}^{H}$. Here, we also use $\mathbf{S}$ to estimate $\mathcal{C}(\mathbf{A})$. Eq. (4) makes problem (3) become the classic MLFA estimation problem [2]:
$\min_{\mathbf{S},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}.$ (5)
For solving problem (5), the IMLSE method utilizes a fixed-point iteration and sequentially (not simultaneously) updates the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration by explicit formulas [7]. However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [8]. Hence, they propose the FAAN method with excellent convergence properties. We outline the FAAN method in the next subsection.
III-B FAAN Method
When a direct minimization over all free parameters is very difficult, an iterative (coordinate descent) approach or alternating minimization can always be adopted [14]. As mentioned before, the FAAN method, composed of two steps at each iteration, has been proposed for problem (5). Specifically, since $\mathbf{C}=\mathbf{SS}^{H}+\mathbf{Q}$ makes updating the estimates of $\mathbf{S}$ and $\mathbf{Q}$ difficult, a new parametrization of $\mathbf{C}$ is designed as
$\mathbf{C}=\mathbf{Q}^{1/2}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})\mathbf{Q}^{1/2},$ (6)
where $\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}$ and $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}$ is the eigen-decomposition of $\mathbf{Q}^{-1/2}\mathbf{SS}^{H}\mathbf{Q}^{-1/2}$. Thus, problem (5) is changed to
$\min_{\mathbf{U},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{Q})+\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})+$
$\mathrm{Tr}\big{\{}\mathbf{Q}^{-1/2}\hat{\mathbf{R}}\mathbf{Q}^{-1/2}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}}.$ (7)
According to (III-B), the first step at the $k$th iteration of the FAAN method simultaneously obtains $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ by
$\min_{\mathbf{U},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}}\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})+\mathrm{Tr}\big{\{}\tilde{\mathbf{R}}^{(k)}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}},$ (8)
where $\tilde{\mathbf{R}}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1/2}\hat{\mathbf{R}}[\mathbf{Q}^{(k-1)}]^{-1/2}$. This leads to
$\mathbf{U}^{(k)}$ $=$ $\mathbf{B}^{(k)}\in\mathbb{C}^{N\times M},$ (9a)
$\Lambda_{m,m}^{(k)}$ $=$ $(\lambda_{m}^{(k)}-1)_{+},m=1,\dots,M$ (9b)
where the $\lambda_{m}^{(k)}$’s are the eigenvalues of $\tilde{\mathbf{R}}^{(k)}$ and $\lambda_{1}^{(k)}\geq\cdots\geq\lambda_{N}^{(k)}$, $\mathbf{B}^{(k)}$ is composed of the $M$ principal eigenvectors. Moreover, the operator $(\cdot)_{+}$ replaces the negative value with zero and leaves the nonnegative value unchanged.
The second step at the $k$th iteration of the FAAN method sequentially (not simultaneously) obtains $\sigma_{1}^{(k)},\dots,\sigma_{N}^{(k)}$ by
$\min_{\sigma_{n}>0}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\{\mathbf{Q}^{-1/2}\hat{{\mathbf{R}}}\mathbf{Q}^{-1/2}\boldsymbol{\Gamma}^{(k)}\},$ (10)
where $\boldsymbol{\Gamma}^{(k)}=\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}^{-1}$ [^2] We do not adopt $\boldsymbol{\Gamma}^{(k)}=\mathbf{I}_{N}+\mathbf{U}^{(k)}\big{(}[\boldsymbol{\Lambda}^{(k)}]^{-1}-\mathbf{I}_{M}\big{)}[\mathbf{U}^{(k)}]^{H}$ since numerical results show its low accuracy.. This leads to
$\sigma_{n}=\frac{1}{2}\Big{(}b_{n}^{(k)}+\sqrt{[b_{n}^{(k)}]^{2}+4c_{n}^{(k)}}\Big{)}>0,n=1,\dots,N$ (11)
where $\Re(\cdot)$ denotes the real part of the complex value, $b_{n}^{(k)}=\sum_{i\neq n}\Re\{\hat{R}_{i,n}\Gamma_{i,n}^{(k)}\}/\sigma_{i}$ and $c_{n}^{(k)}=\hat{R}_{n,n}\Gamma_{n,n}^{(k)}>0$. The details of the FAAN method are in Algorithm 1.
Algorithm 1 FAAN Method
1: Initialize $\mathbf{Q}^{(0)}$, $K_{1}$, $K_{2}$, and $k=1$.
2:while $k\leq K_{1}$ do
3:  Obtain $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ using (9).
4:  Obtain $\mathbf{Q}^{(k)}$ using (11) $K_{2}$ times.
5:  $k=k+1$.
6:end while
7: Let $\hat{\mathbf{S}}=[\mathbf{Q}^{(K_{1})}]^{1/2}\mathbf{U}^{(K_{1})}[\boldsymbol{\Lambda}^{(K_{1})}]^{1/2}$ and output $\hat{\mathbf{S}}$.
IV ECME Algorithm
The ECME algorithm has been designed for problem (5) in [6] and yields rapider convergence than the EM algorithm in [15]. However, a CM-step at each iteration utilizes Newton-Raphson to update the estimate of $\mathbf{Q}$ by the actual or negative LLF $f(\mathbf{S},\mathbf{Q})=\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}$ in (5), which leads to that this step is complex and the monotonicity cannot be guaranteed. In this section, we again design the ECME algorithm with excellent convergence.
IV-A Algorithm Procedure
For problem (5), we do not adopt the complete data $(\mathbf{Y},\mathbf{X})$ with $\mathbf{X}=\big{[}\mathbf{x}(1)~\cdots~\mathbf{x}(L)\big{]}$ in [6, 15] and adopt the simpler complete data $(\mathbf{X},\mathbf{V})$ with $\mathbf{V}=\big{[}\mathbf{v}(1)~\cdots~\mathbf{v}(L)\big{]}$ [^3] Numerical results show that the ECME algorithm with $(\mathbf{X},\mathbf{V})$ yields the same convergence as that with $(\mathbf{Y},\mathbf{X})$. However, the ECME algorithm with $(\mathbf{X},\mathbf{V})$ needs fewer computations at each iteration.. Thus, we rewrite the snapshots or measurements $\mathbf{Y}$ in (2) by
$\mathbf{y}(t)=\mathbf{S}\mathbf{x}(t)+\mathbf{v}(t),t=1,\dots,L$ (12)
where $\mathbf{x}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})$ contains $M$ common factors and is uncorrelated with $\mathbf{v}(t)$. Then, the complete-data LLF is
$l(\mathbf{X},\mathbf{V};\mathbf{Q})$ $=$ $\ln p(\mathbf{X})+\ln p(\mathbf{V};\mathbf{Q})$ (13)
$=$ $c-L\Big{(}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\{\hat{\mathbf{R}}_{v}\mathbf{Q}^{-1}\}\Big{)}$
where $c$ is a term unrelated to any free parameters in $(\mathbf{S},\mathbf{Q})$, $\hat{\mathbf{R}}_{v}=(1/L)\sum_{t=1}^{L}\mathbf{v}(t)\mathbf{v}^{H}(t)$. With (13), we can construct the ECME algorithm, whose E-step and two CM-steps at the $k$th iteration are derived below.
IV-A1 First CM-step
To avoid providing $\mathbf{S}^{(0)}$, let this step only obtain $\mathbf{S}^{(k)}$ by minimizing $f(\mathbf{S},\mathbf{Q})$ with respect to $\mathbf{S}$, i.e.,
$\min_{\mathbf{S}}\ln\mathrm{Det}\big{(}\mathbf{SS}^{H}+\mathbf{Q}^{(k-1)}\big{)}+\mathrm{Tr}\big{\{}\hat{\mathbf{R}}(\mathbf{SS}^{H}+\mathbf{Q}^{(k-1)})^{-1}\big{\}}.$ (14)
This problem is very classic in MLFA and its classic solution is [2, 3]
$\mathbf{S}^{(k)}=[\mathbf{Q}^{(k-1)}]^{1/2}\mathbf{B}^{(k)}[\boldsymbol{\Lambda}^{(k)}]^{1/2}\in\mathbb{C}^{N\times M},$ (15)
where $\mathbf{B}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ are from (9).
IV-A2 E-step
Given $\mathbf{Y}$, $\mathbf{S}^{(k)}$, and $\mathbf{Q}^{(k-1)}$, this step finds the expectation of (13) by
$l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ $=$ $\mathrm{E}\big{\{}l(\mathbf{X},\mathbf{V};\mathbf{Q})\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\}$ (16)
$=$ $c-L\Big{(}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\{\hat{\mathbf{R}}_{v}^{(k)}\mathbf{Q}^{-1}\}\Big{)}$
where $\boldsymbol{\Omega}^{(k)}=(\mathbf{S}^{(k)},\mathbf{Q}^{(k-1)})$, ${}\mathrm{E}\{\cdot\}$ and ${}\mathrm{D}\{\cdot\}$ means expectation and covariance, respectively. Moreover,
$\hat{\mathbf{R}}^{(k)}_{v}=\mathrm{E}\big{\{}\hat{\mathbf{R}}_{v}\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}=[\boldsymbol{\delta}^{(k)}]^{H}\hat{\mathbf{R}}\boldsymbol{\delta}^{(k)}+\boldsymbol{\Delta}^{(k)}\succ\mathbf{0}_{N}$
where $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}$, $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$, and $\boldsymbol{\Delta}^{(k)}=\mathbf{Q}^{(k-1)}-\mathbf{Q}^{(k-1)}\boldsymbol{\delta}^{(k)}\succeq\mathbf{0}_{N}$ [^4] We do not adopt $[\mathbf{C}^{(k)}]^{-1}=\mathbf{q}^{(k)}-\mathbf{q}^{(k)}\mathbf{S}^{(k)}(\boldsymbol{\Lambda}^{(k)}+\mathbf{I}_{M})^{-1}[\mathbf{S}^{(k)}]^{H}\mathbf{q}^{(k)}$ with $\mathbf{q}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1}$ since numerical results show its low accuracy.. The conditional distribution of $\mathbf{x}(t)$ can be derived from [16], and
$\mathrm{E}\big{\{}\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}$ $=$ $[\boldsymbol{\delta}^{(k)}]^{H}\mathbf{y}(t),$
$\mathrm{D}\big{\{}\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}$ $=$ $\boldsymbol{\Delta}^{(k)}\succeq\mathbf{0}_{N}.$
IV-A3 Second CM-step
This step only obtains $\mathbf{Q}^{(k)}$ by maximizing $l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ in (16) with respect to $\mathbf{Q}$, i.e.,
$\min_{\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\big{\{}\hat{\mathbf{R}}_{v}^{(k)}\mathbf{Q}^{-1}\}.$ (17)
This problem can be decomposed into the $N$ parallel subproblems:
$\Big{(}\min_{\sigma_{n}^{2}>0}\ln(\sigma_{n}^{2})+[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}/\sigma^{2}_{n}\Big{)}~~n=1,\dots,N.$ (18)
The solutions are
$[\sigma^{2}_{n}]^{(k)}$ $=$ $[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}>0,n=1,\dots,N,$
$\Rightarrow\mathbf{Q}^{(k)}$ $=$ $\mathrm{diag}\big{(}\hat{\mathbf{R}}_{v}^{(k)}\big{)}\succ\mathbf{0}_{N},$ (19)
where $\mathrm{diag}(\cdot)$ is the matrix composed of the diagonal part. The details of the ECME algorithm are in Algorithm 2.
Algorithm 2 ECME Algorithm
1: Initialize $\mathbf{Q}^{(0)}$, $K$, and $k=1$.
2:while $k\leq K$ do
3:  Obtain $\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}$ using (15).
4:  Obtain $\mathbf{Q}^{(k)}$ using (IV-A3).
5:  $k=k+1$.
6:end while
7: Let $\hat{\mathbf{S}}=\mathbf{S}^{(K)}$ and output $\hat{\mathbf{S}}$.
IV-B Convergence and Complexity
From [5, 17], we know that the ECME algorithm satisfies certain regularity conditions and its sequence $\{(\mathbf{S}^{(k)},\mathbf{Q}^{(k)})\}$ stably converges to a stationary point of $f(\mathbf{S},\mathbf{Q})$. Moreover, Eq. (15) and Eq. (IV-A3) guarantee that the solution provided by the ECME algorithm is feasible for problem (5).
By comparing Algorithm 1 and Algorithm 2, we can find that at the $k$th iteration, the computational complexity of Algorithm 1 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (9) and the inverse of $\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}\in\mathbb{C}^{N\times N}$ in (11). Also, the computational complexity of Algorithm 2 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (15) and the inverse of $\mathbf{C}^{(k)}\in\mathbb{C}^{N\times N}$ in (IV-A3). Thus, the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method.
V Numerical Results
Simulation results are used to show and compare Algorithm 1 and Algorithm 2. Let $N=6$, $M=2$, $\mathbf{P}=10\times\mathbf{I}_{M}$, $\mathbf{Q}=\mathrm{diag}\{10,2,3,2,1,3\}$, $\theta_{1}=60^{\circ}$, $\theta_{2}=120^{\circ}$, and $K_{2}=100$. Moreover, $\mathbf{Q}^{(0)}=\mathbf{I}_{N}$.
Without loss of generality, Fig. 1 illustrates the convergence performances of both algorithms given the same snapshots. Here, $L=100$. We can clearly see that the ECME algorithm yields faster convergence than the FAAN method. Because the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method, it is computationally more efficient.
Figure 1: Convergence comparison of the FAAN method and the ECME algorithm.
After obtaining $\hat{\mathbf{S}}$ or the signal subspace estimate $\mathcal{C}(\hat{\mathbf{S}})$, we perform the eigen-decomposition of $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$ to obtain $\hat{\mathbf{U}}_{n}\in\mathbb{C}^{N\times(N-M)}$, which consists of the eigenvectors corresponding to the $N-M$ smallest eigenvalues. To proceed, we adopt the orthogonal projector $\hat{\boldsymbol{\Pi}}_{n}=\hat{\mathbf{U}}_{n}\hat{\mathbf{U}}_{n}^{H}$ [^5] We do not adopt $\hat{\boldsymbol{\Pi}}_{n}=\mathbf{I}_{N}-\hat{\mathbf{S}}(\hat{\mathbf{S}}^{H}\hat{\mathbf{S}})^{-1}\hat{\mathbf{S}}^{H}$ since numerical results show its low accuracy. onto the noise subspace estimate $\mathcal{C}(\hat{\mathbf{U}}_{n})$ and the Root-MUSIC algorithm [18] to obtain the DOA estimates $\hat{\theta}_{m}$’s.
Fig. 2 shows a scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms under 100100100 independent realizations. Here, $L=100$ and $K=K_{1}=1000$. The same snapshots of each realization are processed by both algorithms. We can know that under the same number of iterations $K=K_{1}$, both algorithms provide different $\hat{\mathbf{S}}$’s and for each realization, the $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$ provided by the ECME algorithm is closer to the solution of $\mathbf{S}\mathbf{S}^{H}$ in (5). Hence, Fig. 3 compares the root mean square error (RMSE) performances of DOA estimation related to both algorithms with the Cramer–Rao lower bound (CRLB) [19, 20]. Here, $K=K_{1}=100$ and each RMSE is based on 100010001000 independent realizations. We observe that the RMSEs of DOA estimation from both algorithms approach the CRLBs as $L$ increases. More importantly, since the ECME algorithm yields faster convergence, its $\mathcal{C}(\mathcal{\hat{\mathbf{S}}})$ is more accurate under $K=K_{1}$ and hence its RMSE performance of DOA estimation is better.
Figure 2: Scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from the FAAN method and the ECME algorithm.
Figure 3: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from the FAAN method and the ECME algorithm.
VI Conclusion
In this letter, the ECME algorithm, an extension of the EM algorithm, has been designed for MLFA, which has almost the same computational complexity at each iteration as the FAAN method with excellent convergence properties. However, numerical results have shown that the ECME algorithm yields faster stable convergence and is computationally more efficient. Importantly, the signal subspace estimated by the ECME algorithm can be used for the subspace based DOA estimation in unknown nonuniform noise.

### v2

@@
 The ECME Algorithm Using Factor Analysis for DOA Estimation in Nonuniform Noise
-Abstract
-Maximum likelihood factor analysis has been used for direction of arrival estimation in unknown nonuniform noise and some iterative approaches have been developed.
-In particular, the Factor Analysis for Anisotropic Noise (FAAN) method proposed by Stoica and Babu has excellent convergence properties.
-In this letter, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the expectation-maximization algorithm, is designed, which has almost the same computational complexity at each iteration as the FAAN method.
-However, numerical results show that the ECME algorithm yields faster stable convergence and is computationally more efficient.
-Index Terms:
-Array signal processing, direction of arrival (DOA) estimation, expectation-maximization (EM) algorithm, factor analysis, maximum likelihood.
-I Introduction
+The ECME Algorithm Using Factor Analysis for DOA Estimation in Nonuniform Noise
+Mingyan Gong
+1 Introduction
 Factor analysis (FA) has been a tool in psychology for over one hundred years [1] and its purpose is to extract or estimate the low-rank covariance matrix from a measurement covariance matrix.
-To this end, several techniques have been developed, but the only one available, which provides efficient and consistent estimators, is maximum likelihood FA (MLFA) [2].
+To this end, several techniques have been developed, but the only one available, which provides efficient and consistent estimators, is maximum likelihood FA (MLFA) [2, 3, 4].
 Unfortunately, these estimators cannot be expressed in closed form, which leads to a variety of iterative approaches [5].
@@
 To be specific, each iteration of the EM algorithm consists of an expectation step (E-step) and a maximization step (M-step).
-However, the M-step requires maximizing the complete-data LLF, which causes two shortcomings: 1) slow convergence, and 2) difficult maximization in numerous complex cases.
-To make up for both shortcomings, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the EM algorithm, has been proposed [5] and then applied to MLFA [6].
-However, a conditional M-step (CM-step) at each iteration utilizes Newton-Raphson to update the estimates of partial parameters by the actual LLF and thus is complex as well as unstable.
+However, the M-step requires maximizing the expected complete-data LLF provided by the E-step, which causes two shortcomings: 1) slow convergence, and 2) difficult maximization in numerous complex cases.
+To make up for both shortcomings, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the EM algorithm, has been proposed [7] and later applied to MLFA [8].
+However, a conditional M-step (CM-step) at its each iteration utilizes Newton-Raphson, instead of an explicit formula, to update the estimates of partial parameters by the actual LLF and thus is complex as well as unstable.
 In addition to psychometrics, econometrics, statistics and so on, FA recently finds applications in array signal processing, especially including direction of arrival (DOA) estimation in unknown nonuniform noise [9, 10].
-The nonuniform noise assumption can let the signal received at an array be modeled by FA so that MLFA can be utilized to estimate the signal subspace for the subspace based DOA estimation [9].
+The nonuniform noise assumption can let the signal received at an array be modeled by FA so that MLFA is used to estimate the signal subspace for the subspace based DOA estimation [11].
 For this purpose, an iterative approach, called Iterative ML Subspace Estimation (IMLSE), is proposed in [9].
@@
 However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [10].
-Then, they propose a new iterative approach, called Factor Analysis for Anisotropic Noise (FAAN), with excellent convergence properties.
-In this letter, the ECME algorithm is designed again for MLFA, which has almost the same computational complexity at each iteration as the FAAN method.
+Afterwards, they propose a novel iterative approach called Factor Analysis for Anisotropic Noise (FAAN), which has excellent convergence properties.
+In this article, the ECME algorithm is designed again for MLFA, which has almost the same computational complexity at each iteration as the FAAN method.
 However, numerical results show that the ECME algorithm yields faster stable convergence and is computationally more efficient.
-Importantly, the signal subspace estimated by the ECME algorithm can be used for the subspace based DOA estimation in unknown nonuniform noise.
+Importantly, the signal subspace estimated by the ECME algorithm can be employed for the subspace based DOA estimation in unknown nonuniform noise.
 Notations: $(\cdot)^{T}$ and $(\cdot)^{H}$ mean the transpose and conjugate transpose operators, respectively.
@@
 $(\cdot)^{(k)}$ is an iterative value obtained at the $k$th iteration.
-II Problem Description
+2 Problem Description
 We consider a uniform linear array (ULA) consisting of $N$ isotropic sensors for analytical simplicity.
@@
 $\mathbf{y}(t)=\sum_{m=1}^{M}\mathbf{a}(\theta_{m})s_{m}(t)+\mathbf{v}(t)=\mathbf{A}\mathbf{s}(t)+\mathbf{v}(t),$ (1)
-where $\theta_{m}\in(0,\pi)$ represents the DOA of the $m$th source, $\mathbf{a}(\theta_{m})=[1~\cdots~e^{-(N-1)\jmath\pi\cos(\theta_{m})}]^{T}\in\mathbb{C}^{N}$ is the steering vector of the $m$th source, and $\mathbf{A}=[\mathbf{a}(\theta_{1})~\cdots~\mathbf{a}(\theta_{M})]$ is the array manifold matrix.
-Additionally, $\mathbf{s}(t)=[s_{1}(t)~\cdots~s_{M}(t)]^{T}\in\mathbb{C}^{M}$ is the source signal vector, $\mathbf{v}(t)$ means a spatially nonuniform Gaussian white noise vector with mean zero and covariance $\mathbf{Q}=\mathrm{diag}\{\sigma^{2}_{1},\dots,\sigma^{2}_{N}\}\succ\mathbf{0}_{N}$, i.e., $\mathbf{v}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{Q})$.
+where $\theta_{m}\in(0,\pi)$ represents the DOA of the $m$th source signal, $\mathbf{a}(\theta_{m})=[1~\cdots~e^{-(N-1)\jmath\pi\cos(\theta_{m})}]^{T}\in\mathbb{C}^{N}$ is the steering vector of the $m$th source, and $\mathbf{A}=[\mathbf{a}(\theta_{1})~\cdots~\mathbf{a}(\theta_{M})]$ is the array manifold matrix.
+Additionally, $\mathbf{s}(t)=[s_{1}(t)~\cdots~s_{M}(t)]^{T}\in\mathbb{C}^{M}$ is the source signal vector, $\mathbf{v}(t)$ is a spatially nonuniform Gaussian white noise vector with mean zero and diagonal covariance matrix $\mathbf{Q}=\mathrm{diag}\{\sigma^{2}_{1},\dots,\sigma^{2}_{N}\}\succ\mathbf{0}_{N}$, i.e., $\mathbf{v}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{Q})$.
 When the stochastic source signal model is adopted, $\mathbf{s}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{P})$, where $\mathbf{P}\succeq\mathbf{0}_{M}$ is the source covariance matrix.
@@
 $=$ $-LN\ln(\pi)-L\Big{(}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}\Big{)}$
-where $\boldsymbol{\Phi}=(\boldsymbol{\theta},\mathbf{P},\mathbf{Q})$ with $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$, $p(\cdot)$ is the probability density function, and $\hat{\mathbf{R}}=(1/L)\sum_{t=1}^{L}\mathbf{y}(t)\mathbf{y}^{H}(t)\succ\mathbf{0}_{N}$ [^1] In this letter, we assume $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$, which requires $L\geq N$..
+where $\boldsymbol{\Phi}=(\boldsymbol{\theta},\mathbf{P},\mathbf{Q})$ with $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$, $p(\cdot)$ is the probability density function, and $\hat{\mathbf{R}}=(1/L)\sum_{t=1}^{L}\mathbf{y}(t)\mathbf{y}^{H}(t)\succ\mathbf{0}_{N}$ [^1] In this article, we assume $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$, which requires $L\geq N$..
 Importantly, we need to solve the equivalent ML direction finding problem [13, 14]:
 $\min_{\boldsymbol{\theta}\in(0,\pi)^{M},\mathbf{P}\succeq\mathbf{0}_{M},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}.$ (3)
-In order to solve problem (3), the ECME algorithm has been designed in the previous work [13]. In the next section, we will design the ECME algorithm using FA for the subspace based DOA estimation.
-III MLFA Based Subspace Estimation
+In order to solve problem (3), the ECME algorithm has been designed in the previous work [15]. In next sections, we will design the ECME algorithm using FA for the subspace based DOA estimation.
+3 Subspace Estimation Using MLFA
 The ML direction finding problem (3) is intractable since $\mathbf{C}=\mathbf{APA}^{H}+\mathbf{Q}$ has excessive constraints on free parameters (unknowns) to be estimated.
 To alleviate this issue, we should release as many constraints as possible and introduce as many free parameters as possible.
-Thus, we aim not to solve problem (3) in this letter.
-III-A MLFA Estimation Problem
+Thus, we aim not to solve problem (3) in this article.
+3.1 MLFA Estimation Problem
 It is well-known that the subspace based DOA estimation only requires the signal subspace $\mathcal{C}(\mathbf{A})$ [11]. As a result, let $\mathbf{C}$ be the population covariance matrix in FA [3]
@@
 Eq.
-(4) makes problem (3) become the classic MLFA estimation problem [2]:
+(4) makes problem (3) become the classic MLFA estimation problem [2, 4, 3, 5]:
 $\min_{\mathbf{S},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}.$ (5)
-For solving problem (5), the IMLSE method utilizes a fixed-point iteration and sequentially (not simultaneously) updates the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration by explicit formulas [7].
+For solving problem (5), the IMLSE method utilizes a fixed-point iteration and sequentially (not simultaneously) updates the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration by two explicit formulas [9].
 However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [10].
@@
 We outline the FAAN method in the next subsection.
-III-B FAAN Method
+3.2 FAAN Method
 When a direct minimization over all free parameters is very difficult, an iterative (coordinate descent) approach or alternating minimization can always be adopted [16].
-As mentioned before, the FAAN method, composed of two steps at each iteration, has been proposed for problem (5).
+As mentioned before, the FAAN method has been proposed for problem (5) and its each iteration consists of two steps.
 Specifically, since $\mathbf{C}=\mathbf{SS}^{H}+\mathbf{Q}$ makes updating the estimates of $\mathbf{S}$ and $\mathbf{Q}$ difficult, a new parametrization of $\mathbf{C}$ is designed as
 $\mathbf{C}=\mathbf{Q}^{1/2}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})\mathbf{Q}^{1/2},$ (6)
-where $\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}$ and $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}$ is the eigen-decomposition of $\mathbf{Q}^{-1/2}\mathbf{SS}^{H}\mathbf{Q}^{-1/2}$.
+where $\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}$, $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}$ is the eigen-decomposition of $\mathbf{Q}^{-1/2}\mathbf{SS}^{H}\mathbf{Q}^{-1/2}$, and $\mathbf{S}=\mathbf{Q}^{1/2}\mathbf{U}\boldsymbol{\Lambda}^{1/2}$.
 Thus, problem (5) is changed to
-$\min_{\mathbf{U},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{Q})+\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})+$
-$\mathrm{Tr}\big{\{}\mathbf{Q}^{-1/2}\hat{\mathbf{R}}\mathbf{Q}^{-1/2}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}}.$ (7)
-According to (III-B), the first step at the $k$th iteration of the FAAN method simultaneously obtains $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ by
-$\min_{\mathbf{U},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}}\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})+\mathrm{Tr}\big{\{}\tilde{\mathbf{R}}^{(k)}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}},$ (8)
+$\min_{\mathbf{U}^{H}\mathbf{U}=\mathbf{I}_{M},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{Q})+\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})$
+$+\mathrm{Tr}\big{\{}\mathbf{Q}^{-1/2}\hat{\mathbf{R}}\mathbf{Q}^{-1/2}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}}.$ (7)
+According to (3.2), the first step at the $k$th iteration of the FAAN method simultaneously obtains $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ by
+$\min_{\mathbf{U}^{H}\mathbf{U}=\mathbf{I}_{M},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}}\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})+$
+$\mathrm{Tr}\big{\{}\tilde{\mathbf{R}}^{(k)}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big{\}},$ (8)
 where $\tilde{\mathbf{R}}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1/2}\hat{\mathbf{R}}[\mathbf{Q}^{(k-1)}]^{-1/2}$. This leads to
 $\mathbf{U}^{(k)}$ $=$ $\mathbf{B}^{(k)}\in\mathbb{C}^{N\times M},$ (9a)
-$\Lambda_{m,m}^{(k)}$ $=$ $(\lambda_{m}^{(k)}-1)_{+},m=1,\dots,M$ (9b)
+$\Lambda_{m,m}^{(k)}$ $=$ $(\lambda_{m}^{(k)}-1)_{+}\geq 0,m=1,\dots,M$ (9b)
 where the $\lambda_{m}^{(k)}$’s are the eigenvalues of $\tilde{\mathbf{R}}^{(k)}$ and $\lambda_{1}^{(k)}\geq\cdots\geq\lambda_{N}^{(k)}$, $\mathbf{B}^{(k)}$ is composed of the $M$ principal eigenvectors.
@@
 This leads to
-$\sigma_{n}=\frac{1}{2}\Big{(}b_{n}^{(k)}+\sqrt{[b_{n}^{(k)}]^{2}+4c_{n}^{(k)}}\Big{)}>0,n=1,\dots,N$ (11)
+$\sigma_{n}$ $=$ $\frac{1}{2}\Big{(}b_{n}^{(k)}+\sqrt{[b_{n}^{(k)}]^{2}+4c_{n}^{(k)}}\Big{)}>0,n=1,\dots,N$
 where $\Re(\cdot)$ denotes the real part of the complex value, $b_{n}^{(k)}=\sum_{i\neq n}\Re\{\hat{R}_{i,n}\Gamma_{i,n}^{(k)}\}/\sigma_{i}$ and $c_{n}^{(k)}=\hat{R}_{n,n}\Gamma_{n,n}^{(k)}>0$.
@@
 Algorithm 1 FAAN Method
-1: Initialize $\mathbf{Q}^{(0)}$, $K_{1}$, $K_{2}$, and $k=1$.
+1: Initialize $\mathbf{Q}^{(0)}\succ\mathbf{0}_{N}$, $K_{1}$, $K_{2}$, and $k=1$.
 2:while $k\leq K_{1}$ do
 3:  Obtain $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ using (9).
-4:  Obtain $\mathbf{Q}^{(k)}$ using (11) $K_{2}$ times.
+4:  Obtain $\mathbf{Q}^{(k)}$ using (3.2) $K_{2}$ times.
 5:  $k=k+1$.
@@
 7: Let $\hat{\mathbf{S}}=[\mathbf{Q}^{(K_{1})}]^{1/2}\mathbf{U}^{(K_{1})}[\boldsymbol{\Lambda}^{(K_{1})}]^{1/2}$ and output $\hat{\mathbf{S}}$.
-IV ECME Algorithm
+4 ECME Algorithm
 The ECME algorithm has been designed for problem (5) in [8] and yields rapider convergence than the EM algorithm in [17].
-However, a CM-step at each iteration utilizes Newton-Raphson to update the estimate of $\mathbf{Q}$ by the actual or negative LLF $f(\mathbf{S},\mathbf{Q})=\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}$ in (5), which leads to that this step is complex and the monotonicity cannot be guaranteed.
+However, a CM-step at its each iteration utilizes Newton-Raphson, instead of an explicit formula, to update the estimate of $\mathbf{Q}$ by the actual or negative LLF $f(\mathbf{S},\mathbf{Q})=\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}$ in (5), leading to that this step is complex and the monotonicity cannot be guaranteed.
 In this section, we again design the ECME algorithm with excellent convergence.
-IV-A Algorithm Procedure
+4.1 Algorithm Procedure
 For problem (5), we do not adopt the complete data $(\mathbf{Y},\mathbf{X})$ with $\mathbf{X}=\big{[}\mathbf{x}(1)~\cdots~\mathbf{x}(L)\big{]}$ in [8, 17] and adopt the simpler complete data $(\mathbf{X},\mathbf{V})$ with $\mathbf{V}=\big{[}\mathbf{v}(1)~\cdots~\mathbf{v}(L)\big{]}$ [^3] Numerical results show that the ECME algorithm with $(\mathbf{X},\mathbf{V})$ yields the same convergence as that with $(\mathbf{Y},\mathbf{X})$.
 However, the ECME algorithm with $(\mathbf{X},\mathbf{V})$ needs fewer computations at each iteration..
-Thus, we rewrite the snapshots or measurements $\mathbf{Y}$ in (2) by
+Thus, we rewrite the snapshots or measurements $\mathbf{Y}$ in (2) by [2, 4, 3]
 $\mathbf{y}(t)=\mathbf{S}\mathbf{x}(t)+\mathbf{v}(t),t=1,\dots,L$ (12)
-where $\mathbf{x}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})$ contains $M$ common factors and is uncorrelated with $\mathbf{v}(t)$. Then, the complete-data LLF is
-$l(\mathbf{X},\mathbf{V};\mathbf{Q})$ $=$ $\ln p(\mathbf{X})+\ln p(\mathbf{V};\mathbf{Q})$ (13)
+where $\mathbf{x}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})$ is composed of $M$ common factors and uncorrelated with $\mathbf{v}(t)$. Then, the complete-data LLF is
+$l(\mathbf{X},\mathbf{V};\mathbf{Q})$ $=$ $\ln p(\mathbf{X})+\ln p(\mathbf{V};\mathbf{Q})$
 $=$ $c-L\Big{(}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\{\hat{\mathbf{R}}_{v}\mathbf{Q}^{-1}\}\Big{)}$
 where $c$ is a term unrelated to any free parameters in $(\mathbf{S},\mathbf{Q})$, $\hat{\mathbf{R}}_{v}=(1/L)\sum_{t=1}^{L}\mathbf{v}(t)\mathbf{v}^{H}(t)$.
-With (13), we can construct the ECME algorithm, whose E-step and two CM-steps at the $k$th iteration are derived below.
-IV-A1 First CM-step
-To avoid providing $\mathbf{S}^{(0)}$, let this step only obtain $\mathbf{S}^{(k)}$ by minimizing $f(\mathbf{S},\mathbf{Q})$ with respect to $\mathbf{S}$, i.e.,
+With (4.1), we can construct the ECME algorithm, whose E-step and two CM-steps at the $k$th iteration are derived below.
+4.1.1 First CM-step:
+Given $\mathbf{Q}=\mathbf{Q}^{(k-1)}$, let this step only obtain $\mathbf{S}^{(k)}$ by minimizing the actual or negative LLF $f(\mathbf{S},\mathbf{Q})$ with respect to $\mathbf{S}$, i.e.,
 $\min_{\mathbf{S}}\ln\mathrm{Det}\big{(}\mathbf{SS}^{H}+\mathbf{Q}^{(k-1)}\big{)}+\mathrm{Tr}\big{\{}\hat{\mathbf{R}}(\mathbf{SS}^{H}+\mathbf{Q}^{(k-1)})^{-1}\big{\}}.$ (14)
-This problem is very classic in MLFA and its classic solution is [2, 3]
-$\mathbf{S}^{(k)}=[\mathbf{Q}^{(k-1)}]^{1/2}\mathbf{B}^{(k)}[\boldsymbol{\Lambda}^{(k)}]^{1/2}\in\mathbb{C}^{N\times M},$ (15)
-where $\mathbf{B}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ are from (9).
-IV-A2 E-step
-Given $\mathbf{Y}$, $\mathbf{S}^{(k)}$, and $\mathbf{Q}^{(k-1)}$, this step finds the expectation of (13) by
-$l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ $=$ $\mathrm{E}\big{\{}l(\mathbf{X},\mathbf{V};\mathbf{Q})\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\}$ (16)
+This problem is classic [2, 4, 3] but the classic solutions in [18, 19] may not exist. Thus, we resort to the FAAN method by
+$\mathbf{S}^{(k)}=[\mathbf{Q}^{(k-1)}]^{1/2}\mathbf{U}^{(k)}[\boldsymbol{\Lambda}^{(k)}]^{1/2}\in\mathbb{C}^{N\times M},$ (15)
+where $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ are from (9).
+4.1.2 E-step:
+Given $\mathbf{Y}$, $\mathbf{S}^{(k)}$, and $\mathbf{Q}^{(k-1)}$, this step finds the expectation of $l(\mathbf{X},\mathbf{V};\mathbf{Q})$ in (4.1) by
+$l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ $=$ $\mathrm{E}\big{\{}l(\mathbf{X},\mathbf{V};\mathbf{Q})\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\}$
 $=$ $c-L\Big{(}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\{\hat{\mathbf{R}}_{v}^{(k)}\mathbf{Q}^{-1}\}\Big{)}$
@@
 $\hat{\mathbf{R}}^{(k)}_{v}=\mathrm{E}\big{\{}\hat{\mathbf{R}}_{v}\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}=[\boldsymbol{\delta}^{(k)}]^{H}\hat{\mathbf{R}}\boldsymbol{\delta}^{(k)}+\boldsymbol{\Delta}^{(k)}\succ\mathbf{0}_{N}$
-where $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}$, $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$, and $\boldsymbol{\Delta}^{(k)}=\mathbf{Q}^{(k-1)}-\mathbf{Q}^{(k-1)}\boldsymbol{\delta}^{(k)}\succeq\mathbf{0}_{N}$ [^4] We do not adopt $[\mathbf{C}^{(k)}]^{-1}=\mathbf{q}^{(k)}-\mathbf{q}^{(k)}\mathbf{S}^{(k)}(\boldsymbol{\Lambda}^{(k)}+\mathbf{I}_{M})^{-1}[\mathbf{S}^{(k)}]^{H}\mathbf{q}^{(k)}$ with $\mathbf{q}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1}$ since numerical results show its low accuracy..
-The conditional distribution of $\mathbf{x}(t)$ can be derived from [16], and
+where $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}$, $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$ [^4] We do not adopt $\mathbf{C}^{(k)}]^{-1}=\mathbf{q}^{(k)}-\mathbf{q}^{(k)}\mathbf{S}^{(k)}(\boldsymbol{\Lambda}^{(k)}+\mathbf{I}_{M})^{-1}[\mathbf{S}^{(k)}]^{H}\mathbf{q}^{(k)}$ with $\mathbf{q}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1}$ since numerical results show its low accuracy., and $\boldsymbol{\Delta}^{(k)}=\mathbf{Q}^{(k-1)}-\mathbf{Q}^{(k-1)}\boldsymbol{\delta}^{(k)}\succeq\mathbf{0}_{N}$.
+The conditional distribution of $\mathbf{v}(t)$ can be derived from [20], and
 $\mathrm{E}\big{\{}\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}$ $=$ $[\boldsymbol{\delta}^{(k)}]^{H}\mathbf{y}(t),$
 $\mathrm{D}\big{\{}\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}$ $=$ $\boldsymbol{\Delta}^{(k)}\succeq\mathbf{0}_{N}.$
-IV-A3 Second CM-step
-This step only obtains $\mathbf{Q}^{(k)}$ by maximizing $l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ in (16) with respect to $\mathbf{Q}$, i.e.,
+4.1.3 Second CM-step:
+This step only obtains $\mathbf{Q}^{(k)}$ by maximizing $l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ in (4.1.2) with respect to $\mathbf{Q}$, i.e.,
 $\min_{\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\big{\{}\hat{\mathbf{R}}_{v}^{(k)}\mathbf{Q}^{-1}\}.$ (17)
@@
 Algorithm 2 ECME Algorithm
-1: Initialize $\mathbf{Q}^{(0)}$, $K$, and $k=1$.
+1: Initialize $\mathbf{Q}^{(0)}\succ\mathbf{0}_{N}$, $K$, and $k=1$.
 2:while $k\leq K$ do
 3:  Obtain $\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}$ using (15).
-4:  Obtain $\mathbf{Q}^{(k)}$ using (IV-A3).
+4:  Obtain $\mathbf{Q}^{(k)}$ using (4.1.3).
 5:  $k=k+1$.
@@
 7: Let $\hat{\mathbf{S}}=\mathbf{S}^{(K)}$ and output $\hat{\mathbf{S}}$.
-IV-B Convergence and Complexity
-From [5, 17], we know that the ECME algorithm satisfies certain regularity conditions and its sequence $\{(\mathbf{S}^{(k)},\mathbf{Q}^{(k)})\}$ stably converges to a stationary point of $f(\mathbf{S},\mathbf{Q})$.
+4.2 Convergence and Complexity
+From [7, 21], we know that the ECME algorithm satisfies certain regularity conditions and its sequence $\big{\{}\big{(}\mathbf{S}^{(k)},\mathbf{Q}^{(k)}\big{)}\big{\}}$ always converges to a stationary point of $f(\mathbf{S},\mathbf{Q})$.
+Fortunately, simulation results in [10] and the subsequent section show that although the objective function $f(\mathbf{S},\mathbf{Q)}$ in problem (5) is non-convex, it seems to have a favourable landscape, which makes its minimization by the FAAN method or the ECME algorithm a relatively straightforward task.
 Moreover, Eq.
 (15) and Eq.
-(IV-A3) guarantee that the solution provided by the ECME algorithm is feasible for problem (5).
-By comparing Algorithm 1 and Algorithm 2, we can find that at the $k$th iteration, the computational complexity of Algorithm 1 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (9) and the inverse of $\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}\in\mathbb{C}^{N\times N}$ in (11).
-Also, the computational complexity of Algorithm 2 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (15) and the inverse of $\mathbf{C}^{(k)}\in\mathbb{C}^{N\times N}$ in (IV-A3).
+(4.1.3) guarantee that the solution provided by the ECME algorithm is “proper" for problem (5) [18, 22].
+By comparing Algorithm 1 and Algorithm 2, we can find that at the $k$th iteration, the computational complexity of Algorithm 1 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (9) and the inverse of $\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}\in\mathbb{C}^{N\times N}$ in (3.2).
+Also, the computational complexity of Algorithm 2 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (15) and the inverse of $\mathbf{C}^{(k)}\in\mathbb{C}^{N\times N}$ in (4.1.3).
 Thus, the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method.
-V Numerical Results
+5 Numerical Results
 Simulation results are used to show and compare Algorithm 1 and Algorithm 2.
@@
 Figure 1: Convergence comparison of the FAAN method and the ECME algorithm.
-After obtaining $\hat{\mathbf{S}}$ or the signal subspace estimate $\mathcal{C}(\hat{\mathbf{S}})$, we perform the eigen-decomposition of $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$ to obtain $\hat{\mathbf{U}}_{n}\in\mathbb{C}^{N\times(N-M)}$, which consists of the eigenvectors corresponding to the $N-M$ smallest eigenvalues.
-To proceed, we adopt the orthogonal projector $\hat{\boldsymbol{\Pi}}_{n}=\hat{\mathbf{U}}_{n}\hat{\mathbf{U}}_{n}^{H}$ [^5] We do not adopt $\hat{\boldsymbol{\Pi}}_{n}=\mathbf{I}_{N}-\hat{\mathbf{S}}(\hat{\mathbf{S}}^{H}\hat{\mathbf{S}})^{-1}\hat{\mathbf{S}}^{H}$ since numerical results show its low accuracy.
-onto the noise subspace estimate $\mathcal{C}(\hat{\mathbf{U}}_{n})$ and the Root-MUSIC algorithm [18] to obtain the DOA estimates $\hat{\theta}_{m}$’s.
-Fig.
-2 shows a scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms under 100100100 independent realizations.
-Here, $L=100$ and $K=K_{1}=1000$.
-The same snapshots of each realization are processed by both algorithms.
-We can know that under the same number of iterations $K=K_{1}$, both algorithms provide different $\hat{\mathbf{S}}$’s and for each realization, the $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$ provided by the ECME algorithm is closer to the solution of $\mathbf{S}\mathbf{S}^{H}$ in (5).
-Hence, Fig.
-3 compares the root mean square error (RMSE) performances of DOA estimation related to both algorithms with the Cramer–Rao lower bound (CRLB) [19, 20].
-Here, $K=K_{1}=100$ and each RMSE is based on 100010001000 independent realizations.
-We observe that the RMSEs of DOA estimation from both algorithms approach the CRLBs as $L$ increases.
-More importantly, since the ECME algorithm yields faster convergence, its $\mathcal{C}(\mathcal{\hat{\mathbf{S}}})$ is more accurate under $K=K_{1}$ and hence its RMSE performance of DOA estimation is better.
+[…]
+[…]
 Figure 2: Scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from the FAAN method and the ECME algorithm.
 Figure 3: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from the FAAN method and the ECME algorithm.
-VI Conclusion
-In this letter, the ECME algorithm, an extension of the EM algorithm, has been designed for MLFA, which has almost the same computational complexity at each iteration as the FAAN method with excellent convergence properties.
+6 Conclusion
+In this article, the ECME algorithm, an extension of the EM algorithm, has been designed for MLFA, which has almost the same computational complexity at each iteration as the FAAN method with excellent convergence properties.
 However, numerical results have shown that the ECME algorithm yields faster stable convergence and is computationally more efficient.
-Importantly, the signal subspace estimated by the ECME algorithm can be used for the subspace based DOA estimation in unknown nonuniform noise.
+Importantly, the signal subspace estimated by the ECME algorithm can be employed for the subspace based DOA estimation in unknown nonuniform noise.

### v3

@@
 1 Introduction
-Factor analysis (FA) has been a tool in psychology for over one hundred years [1] and its purpose is to extract or estimate the low-rank covariance matrix from a measurement covariance matrix.
-To this end, several techniques have been developed, but the only one available, which provides efficient and consistent estimators, is maximum likelihood FA (MLFA) [2, 3, 4].
-Unfortunately, these estimators cannot be expressed in closed form, which leads to a variety of iterative approaches [5].
-In particular, one of the most promising is the Expectation-Maximization (EM) algorithm [6].
-The EM algorithm regards observations or measurements as the incomplete data and designs the complete data containing missing (latent or unobserved) data.
+Factor analysis (FA) has been a tool in psychology for over one hundred years [spearman] and its purpose is to extract or estimate the low-rank covariance matrix from a measurement covariance matrix.
+To this end, several techniques have been developed, but the only one available, which provides an efficient and consistent estimator, is maximum likelihood FA (MLFA) [lawley, basilevsky, mulaik].
+Unfortunately, this ML based estimator cannot be expressed in closed form, which leads to a variety of iterative approaches [jöreskog].
+In particular, one of the most promising is the Expectation-Maximization (EM) algorithm [dempster, rubin].
+The EM algorithm regards observations or measurements as the incomplete data and introduces the complete data containing missing (latent or unobserved) data.
 The complete data is so manipulable that its log-likelihood function (LLF) is more tractable than the incomplete-data or actual LLF.
-Hence, the EM algorithm uses the complete-data LLF to efficiently update the estimates of parameters, which can still increase the actual likelihood.
+Accordingly, the EM algorithm only uses the complete-data LLF to efficiently update the estimates of parameters at each iteration, which can still increase the actual likelihood.
 To be specific, each iteration of the EM algorithm consists of an expectation step (E-step) and a maximization step (M-step).
 However, the M-step requires maximizing the expected complete-data LLF provided by the E-step, which causes two shortcomings: 1) slow convergence, and 2) difficult maximization in numerous complex cases.
-To make up for both shortcomings, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the EM algorithm, has been proposed [7] and later applied to MLFA [8].
+To make up for both shortcomings, the Expectation/Conditional Maximization Either (ECME) algorithm, an extension of the EM algorithm, has been proposed [liu] and later applied to MLFA [liu2].
 However, a conditional M-step (CM-step) at its each iteration utilizes Newton-Raphson, instead of an explicit formula, to update the estimates of partial parameters by the actual LLF and thus is complex as well as unstable.
-In addition to psychometrics, econometrics, statistics and so on, FA recently finds applications in array signal processing, especially including direction of arrival (DOA) estimation in unknown nonuniform noise [9, 10].
-The nonuniform noise assumption can let the signal received at an array be modeled by FA so that MLFA is used to estimate the signal subspace for the subspace based DOA estimation [11].
-For this purpose, an iterative approach, called Iterative ML Subspace Estimation (IMLSE), is proposed in [9].
-This method sequentially updates the estimates of parameters at each iteration by explicit formulas and is computationally attractive.
-However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [10].
-Afterwards, they propose a novel iterative approach called Factor Analysis for Anisotropic Noise (FAAN), which has excellent convergence properties.
-In this article, the ECME algorithm is designed again for MLFA, which has almost the same computational complexity at each iteration as the FAAN method.
-However, numerical results show that the ECME algorithm yields faster stable convergence and is computationally more efficient.
-Importantly, the signal subspace estimated by the ECME algorithm can be employed for the subspace based DOA estimation in unknown nonuniform noise.
+In addition to psychometrics, econometrics, statistics and so on, FA recently finds applications in array signal processing, especially including direction of arrival (DOA) estimation in unknown nonuniform noise [liao, stoica].
+The nonuniform noise assumption can let the signal received at an array be modeled by FA so that MLFA is employed to estimate the signal subspace for the subspace based DOA estimation [krim].
+For this purpose, the iterative approach called Iterative ML Subspace Estimation (IMLSE) is proposed in [liao].
+This method sequentially updates the estimates of parameters at each iteration by two explicit formulas and is computationally attractive.
+However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [stoica].
+Afterwards, they propose the novel iterative approach called Factor Analysis for Anisotropic Noise (FAAN), which has excellent convergence properties.
+In this article, the ECME algorithm is designed again for MLFA by introducing new complete data, which can thus use two explicit formulas to sequentially update the estimates of parameters at each iteration and have excellent convergence properties.
+Theoretical analysis shows that the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method.
+However, numerical results show that the ECME algorithm yields faster stable convergence and the convergence to the global optimum is easier.
+Importantly, MLFA is not the best choice for the subspace based DOA estimation in unknown nonuniform noise.
 Notations: $(\cdot)^{T}$ and $(\cdot)^{H}$ mean the transpose and conjugate transpose operators, respectively.
-$\mathrm{Det}(\boldsymbol{\Sigma})$, $\mathrm{Tr}\{\boldsymbol{\Sigma}\}$, $\boldsymbol{\Sigma}^{1/2}$, and $\boldsymbol{\Sigma}^{-1}$ stands for the determinant, trace, square root, and inverse of square matrix $\boldsymbol{\Sigma}$.
+$\mathrm{Det}(\boldsymbol{\Sigma})$, $\mathrm{Tr}\{\boldsymbol{\Sigma}\}$, $\boldsymbol{\Sigma}^{1/2}$, and $\boldsymbol{\Sigma}^{-1}$ stand for the determinant, trace, square root, and inverse of square matrix $\boldsymbol{\Sigma}$.
 $\Sigma_{i,j}$ is the element in the $i$th row and $j$th column of $\boldsymbol{\Sigma}$.
@@
 $\mathcal{C}(\boldsymbol{\Sigma})$ denotes the column space of $\boldsymbol{\Sigma}$.
-$(\cdot)^{(k)}$ is an iterative value obtained at the $k$th iteration.
+$\theta^{(k)}$ is the iterative value of $\theta$ obtained at the $k$th iteration of an algorithm.
 2 Problem Description
 We consider a uniform linear array (ULA) consisting of $N$ isotropic sensors for analytical simplicity.
-$M<N$ narrow-band source signals are assumed to impinge upon the array from far field and the distance between arbitrary two adjacent sensors is the half-wavelength of the source signals.
+M(M<N)M~(M<N) narrow-band source signals are assumed to impinge upon the array from far field and the distance between arbitrary two adjacent sensors is the half-wavelength of the source signals.
 Thus, the composite signal received at this ULA can be expressed as
 $\mathbf{y}(t)=\sum_{m=1}^{M}\mathbf{a}(\theta_{m})s_{m}(t)+\mathbf{v}(t)=\mathbf{A}\mathbf{s}(t)+\mathbf{v}(t),$ (1)
-where $\theta_{m}\in(0,\pi)$ represents the DOA of the $m$th source signal, $\mathbf{a}(\theta_{m})=[1~\cdots~e^{-(N-1)\jmath\pi\cos(\theta_{m})}]^{T}\in\mathbb{C}^{N}$ is the steering vector of the $m$th source, and $\mathbf{A}=[\mathbf{a}(\theta_{1})~\cdots~\mathbf{a}(\theta_{M})]$ is the array manifold matrix.
-Additionally, $\mathbf{s}(t)=[s_{1}(t)~\cdots~s_{M}(t)]^{T}\in\mathbb{C}^{M}$ is the source signal vector, $\mathbf{v}(t)$ is a spatially nonuniform Gaussian white noise vector with mean zero and diagonal covariance matrix $\mathbf{Q}=\mathrm{diag}\{\sigma^{2}_{1},\dots,\sigma^{2}_{N}\}\succ\mathbf{0}_{N}$, i.e., $\mathbf{v}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{Q})$.
+where $\theta_{m}\in(0,\pi)~(\mathrm{radian})$ represents the DOA of the $m$th source signal, $\mathbf{a}(\theta_{m})=[1~e^{-\jmath\pi\cos(\theta_{m})}~\cdots~e^{-(N-1)\jmath\pi\cos(\theta_{m})}]^{T}\in\mathbb{C}^{N}$ is the steering vector of the $m$th source signal, and $\mathbf{A}=[\mathbf{a}(\theta_{1})~\cdots~\mathbf{a}(\theta_{M})]$ is the array manifold matrix.
+For convenience, let $\theta_{1}<\cdots<\theta_{M}$.
+Additionally, $\mathbf{s}(t)=[s_{1}(t)~\cdots~s_{M}(t)]^{T}\in\mathbb{C}^{M}$ is the source signal vector, $\mathbf{v}(t)$ is the spatially nonuniform Gaussian white noise vector with mean zero and diagonal covariance matrix $\mathbf{Q}=\mathrm{diag}\{\sigma^{2}_{1},\dots,\sigma^{2}_{N}\}\succ\mathbf{0}_{N}$, i.e., $\mathbf{v}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{Q})$.
 When the stochastic source signal model is adopted, $\mathbf{s}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{P})$, where $\mathbf{P}\succeq\mathbf{0}_{M}$ is the source covariance matrix.
-In (1), $\mathbf{s}(t)$ and $\mathbf{v}(t)$ are assumed to be mutually uncorrelated, which leads to $\mathbf{y}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{C})$ with $\mathbf{C}=\mathbf{A}\mathbf{P}\mathbf{A}^{H}+\mathbf{Q}$ [11].
+In (1), $\mathbf{s}(t)$ and $\mathbf{v}(t)$ are assumed to be mutually uncorrelated, which leads to $\mathbf{y}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{C})$ with $\mathbf{C}=\mathbf{A}\mathbf{P}\mathbf{A}^{H}+\mathbf{Q}\succ\mathbf{0}_{N}$ [krim].
 In practice, we first sample $\mathbf{y}(t)$ and obtain independent and identically distributed snapshots as follows:
@@
 where $L$ is the number of snapshots. Then, we obtain the DOA estimates $\hat{\theta}_{m}$’s by processing the snapshots $\mathbf{y}(t)$’s.
-From the perspective of ML direction finding [12], we need to write the LLF of $\mathbf{Y}$ with $\mathbf{Y}=[\mathbf{y}(1)~\cdots~\mathbf{y}(L)]$ as
-$l(\boldsymbol{\Phi})$ $=$ $\ln p(\mathbf{Y};\boldsymbol{\Phi})={\sum}_{t=1}^{L}\ln p\big{(}\mathbf{y}(t);\boldsymbol{\Phi}\big{)}$
-$=$ $-LN\ln(\pi)-L\Big{(}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}\Big{)}$
-where $\boldsymbol{\Phi}=(\boldsymbol{\theta},\mathbf{P},\mathbf{Q})$ with $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$, $p(\cdot)$ is the probability density function, and $\hat{\mathbf{R}}=(1/L)\sum_{t=1}^{L}\mathbf{y}(t)\mathbf{y}^{H}(t)\succ\mathbf{0}_{N}$ [^1] In this article, we assume $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$, which requires $L\geq N$..
-Importantly, we need to solve the equivalent ML direction finding problem [13, 14]:
+From the processing perspective of ML direction finding [kay], we need to write the LLF of $\mathbf{Y}$ with $\mathbf{Y}=[\mathbf{y}(1)~\cdots~\mathbf{y}(L)]$ as
+$l(\boldsymbol{\Phi})$ $=$ $\ln p(\mathbf{Y};\boldsymbol{\Phi})={\sum}_{t=1}^{L}\ln p\big(\mathbf{y}(t);\boldsymbol{\Phi}\big)$ (3)
+$=$ $a-L\Big(\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}\Big)$
+where $a$ is a constant, $\boldsymbol{\Phi}=(\boldsymbol{\theta},\mathbf{P},\mathbf{Q})$ with $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$, $p(\cdot)$ is the probability density function (PDF), and $\hat{\mathbf{R}}=(1/L)\sum_{t=1}^{L}\mathbf{y}(t)\mathbf{y}^{H}(t)\succ\mathbf{0}_{N}$ [^1] In this article, we assume $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$, which requires $L\geq N$.
+is the snapshot covariance matrix.
+Importantly, we need to solve the equivalent ML direction finding problem [jaffer, stoica2, pesavento]:
 $\min_{\boldsymbol{\theta}\in(0,\pi)^{M},\mathbf{P}\succeq\mathbf{0}_{M},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}.$ (4)
-In order to solve problem (3), the ECME algorithm has been designed in the previous work [15]. In next sections, we will design the ECME algorithm using FA for the subspace based DOA estimation.
 3 Subspace Estimation Using MLFA
@@
 3.1 MLFA Estimation Problem
-It is well-known that the subspace based DOA estimation only requires the signal subspace $\mathcal{C}(\mathbf{A})$ [11]. As a result, let $\mathbf{C}$ be the population covariance matrix in FA [3]
+It is well-known that the subspace based DOA estimation only requires the signal subspace $\mathcal{C}(\mathbf{A})$ [krim].
+As a result, let $\mathbf{C}$ be the population covariance matrix in FA [basilevsky]
 $\mathbf{C}=\mathbf{S}\mathbf{S}^{H}+\mathbf{Q},$ (5)
-where all elements in $\mathbf{S}\in\mathbb{C}^{N\times M}$ are free parameters (factor loadings) and hence $\mathbf{SS}^{H}$ must be able to equal any $\mathbf{APA}^{H}$.
-Here, we also use $\mathbf{S}$ to estimate $\mathcal{C}(\mathbf{A})$.
+where all the $NM$ elements in $\mathbf{S}\in\mathbb{C}^{N\times M}$ are free parameters (factor loadings) and hence $\mathbf{SS}^{H}$ must be able to equal any $\mathbf{APA}^{H}$.
+In this article, we also use $\mathbf{S}$ to estimate $\mathcal{C}(\mathbf{A})$.
 Eq.
-(4) makes problem (3) become the classic MLFA estimation problem [2, 4, 3, 5]:
+(5) makes problem (4) become the classic MLFA estimation problem [lawley, basilevsky, mulaik, jöreskog]:
 $\min_{\mathbf{S},\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}.$ (6)
-For solving problem (5), the IMLSE method utilizes a fixed-point iteration and sequentially (not simultaneously) updates the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration by two explicit formulas [9].
-However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [10].
+For solving problem (6), the IMLSE method utilizes a fixed-point iteration and sequentially (not simultaneously) updates the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration by two explicit formulas [liao].
+However, Stoica and Babu have shown that the IMLSE method cannot guarantee its monotonicity and may fail to converge [stoica].
 Hence, they propose the FAAN method with excellent convergence properties.
@@
 3.2 FAAN Method
-When a direct minimization over all free parameters is very difficult, an iterative (coordinate descent) approach or alternating minimization can always be adopted [16].
+When a direct minimization over all free parameters is very difficult, an iterative (coordinate descent) approach or alternating minimization can always be adopted [bezdek].
 As mentioned before, the FAAN method has been proposed for problem (6) and its each iteration consists of two steps.
@@
 $\mathbf{C}=\mathbf{Q}^{1/2}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})\mathbf{Q}^{1/2},$ (7)
-where $\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}$, $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}$ is the eigen-decomposition of $\mathbf{Q}^{-1/2}\mathbf{SS}^{H}\mathbf{Q}^{-1/2}$, and $\mathbf{S}=\mathbf{Q}^{1/2}\mathbf{U}\boldsymbol{\Lambda}^{1/2}$.
+where $\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}$, $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}$ is the eigen-decomposition of $\mathbf{Q}^{-1/2}\mathbf{SS}^{H}\mathbf{Q}^{-1/2}$, and $\mathbf{S}=\mathbf{Q}^{1/2}\mathbf{U}\boldsymbol{\Lambda}^{1/2}$ for convenience.
 Thus, problem (6) is changed to
@@
 $+\mathrm{Tr}\big\{\mathbf{Q}^{-1/2}\hat{\mathbf{R}}\mathbf{Q}^{-1/2}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big\}.$ (8)
-According to (3.2), the first step at the $k$th iteration of the FAAN method simultaneously obtains $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ by
+According to (3.2), given $\mathbf{Q}=\mathbf{Q}^{(k-1)}\succ\mathbf{0}_{N}$, the first step at the $k$th iteration of the FAAN method simultaneously obtains $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ by
 $\min_{\mathbf{U}^{H}\mathbf{U}=\mathbf{I}_{M},\boldsymbol{\Lambda}\succeq\mathbf{0}_{M}}\ln\mathrm{Det}(\boldsymbol{\Lambda}+\mathbf{I}_{M})+$
 $\mathrm{Tr}\big\{\tilde{\mathbf{R}}^{(k)}(\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{H}+\mathbf{I}_{N})^{-1}\big\},$ (9)
-where $\tilde{\mathbf{R}}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1/2}\hat{\mathbf{R}}[\mathbf{Q}^{(k-1)}]^{-1/2}$. This leads to
+where $\tilde{\mathbf{R}}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1/2}\hat{\mathbf{R}}[\mathbf{Q}^{(k-1)}]^{-1/2}\succ\mathbf{0}_{N}$. This leads to
 $\mathbf{U}^{(k)}$ $=$ $\mathbf{B}^{(k)}\in\mathbb{C}^{N\times M},$ (10a)
 $\Lambda_{m,m}^{(k)}$ $=$ $(\lambda_{m}^{(k)}-1)_{+}\geq 0,m=1,\dots,M$ (10b)
-where the $\lambda_{m}^{(k)}$’s are the eigenvalues of $\tilde{\mathbf{R}}^{(k)}$ and $\lambda_{1}^{(k)}\geq\cdots\geq\lambda_{N}^{(k)}$, $\mathbf{B}^{(k)}$ is composed of the $M$ principal eigenvectors.
-Moreover, the operator $(\cdot)_{+}$ replaces the negative value with zero and leaves the nonnegative value unchanged.
-The second step at the $k$th iteration of the FAAN method sequentially (not simultaneously) obtains $\sigma_{1}^{(k)},\dots,\sigma_{N}^{(k)}$ by
+where the $\lambda_{m}^{(k)}$’s are the eigenvalues of $\tilde{\mathbf{R}}^{(k)}$ and $\lambda_{1}^{(k)}\geq\cdots\geq\lambda_{N}^{(k)}>0$, the $m$th column of $\mathbf{B}^{(k)}$ is the eigenvector corresponding to $\lambda_{m}^{(k)}$.
+Moreover, the operator $(b)_{+}$ replaces the negative value of real number $b$ with zero and leaves the nonnegative value unchanged.
+The second step at the $k$th iteration of the FAAN method sequentially (one by one) obtains $\sigma_{1}^{(k)},\dots,\sigma_{N}^{(k)}$ by the $N$ problems:
 $\min_{\sigma_{n}>0}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\{\mathbf{Q}^{-1/2}\hat{{\mathbf{R}}}\mathbf{Q}^{-1/2}\boldsymbol{\Gamma}^{(k)}\},$ (11)
-where $\boldsymbol{\Gamma}^{(k)}=\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}^{-1}$ [^2] We do not adopt $\boldsymbol{\Gamma}^{(k)}=\mathbf{I}_{N}+\mathbf{U}^{(k)}\big{(}[\boldsymbol{\Lambda}^{(k)}]^{-1}-\mathbf{I}_{M}\big{)}[\mathbf{U}^{(k)}]^{H}$ since numerical results show its low accuracy..
+where $n=1,\dots,N$ and $\boldsymbol{\Gamma}^{(k)}=\big(\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big)^{-1}\succ\mathbf{0}_{N}$ [^2] We do not adopt $\boldsymbol{\Gamma}^{(k)}=\mathbf{I}_{N}+\mathbf{U}^{(k)}\big([\boldsymbol{\Lambda}^{(k)}]^{-1}-\mathbf{I}_{M}\big)[\mathbf{U}^{(k)}]^{H}$ since numerical results show its low accuracy..
 This leads to
-$\sigma_{n}$ $=$ $\frac{1}{2}\Big{(}b_{n}^{(k)}+\sqrt{[b_{n}^{(k)}]^{2}+4c_{n}^{(k)}}\Big{)}>0,n=1,\dots,N$
-where $\Re(\cdot)$ denotes the real part of the complex value, $b_{n}^{(k)}=\sum_{i\neq n}\Re\{\hat{R}_{i,n}\Gamma_{i,n}^{(k)}\}/\sigma_{i}$ and $c_{n}^{(k)}=\hat{R}_{n,n}\Gamma_{n,n}^{(k)}>0$.
+$\sigma_{n}=\frac{1}{2}\Big(b_{n}^{(k)}+\sqrt{[b_{n}^{(k)}]^{2}+4c_{n}^{(k)}}\Big)>0$ (12)
+where $n=1,\dots,N$, $\Re(b)$ denotes the real part of complex number $b$, $b_{n}^{(k)}=\sum_{i\neq n}\Re\{\hat{R}_{i,n}\Gamma_{i,n}^{(k)}\}/\sigma_{i}$ and $c_{n}^{(k)}=\hat{R}_{n,n}\Gamma_{n,n}^{(k)}>0$.
 The details of the FAAN method are in Algorithm 1.
@@
 3:  Obtain $\mathbf{U}^{(k)}$ and $\boldsymbol{\Lambda}^{(k)}$ using (10).
-4:  Obtain $\mathbf{Q}^{(k)}$ using (3.2) $K_{2}$ times.
+4:  Obtain $\mathbf{Q}^{(k)}$ using (12) $K_{2}$ times.
 5:  $k=k+1$.
@@
 4 ECME Algorithm
-The ECME algorithm has been designed for problem (5) in [8] and yields rapider convergence than the EM algorithm in [17].
-However, a CM-step at its each iteration utilizes Newton-Raphson, instead of an explicit formula, to update the estimate of $\mathbf{Q}$ by the actual or negative LLF $f(\mathbf{S},\mathbf{Q})=\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}$ in (5), leading to that this step is complex and the monotonicity cannot be guaranteed.
-In this section, we again design the ECME algorithm with excellent convergence.
+The ECME algorithm has been designed for problem (6) in [liu2] and yields rapider convergence than the EM algorithm designed in [rubin].
+However, a CM-step at its each iteration utilizes Newton-Raphson, instead of an explicit formula, to update the estimate of $\mathbf{Q}$ by the objective function or negative LLF $f(\mathbf{S},\mathbf{Q})=\ln\mathrm{Det}(\mathbf{C})+\mathrm{Tr}\{\hat{\mathbf{R}}\mathbf{C}^{-1}\}$ in problem (6), leading to that this step is complex and the monotonicity cannot be guaranteed.
+In this section, we again design the ECME algorithm by introducing new complete data, which can thus use two explicit formulas to sequentially update the estimates of $\mathbf{S}$ and $\mathbf{Q}$ at each iteration and have excellent convergence properties.
 4.1 Algorithm Procedure
-For problem (5), we do not adopt the complete data $(\mathbf{Y},\mathbf{X})$ with $\mathbf{X}=\big{[}\mathbf{x}(1)~\cdots~\mathbf{x}(L)\big{]}$ in [8, 17] and adopt the simpler complete data $(\mathbf{X},\mathbf{V})$ with $\mathbf{V}=\big{[}\mathbf{v}(1)~\cdots~\mathbf{v}(L)\big{]}$ [^3] Numerical results show that the ECME algorithm with $(\mathbf{X},\mathbf{V})$ yields the same convergence as that with $(\mathbf{Y},\mathbf{X})$.
-However, the ECME algorithm with $(\mathbf{X},\mathbf{V})$ needs fewer computations at each iteration..
-Thus, we rewrite the snapshots or measurements $\mathbf{Y}$ in (2) by [2, 4, 3]
+For problem (6), we do not adopt the classic complete data $(\mathbf{Y},\mathbf{X})$ with $\mathbf{X}=\big[\mathbf{x}(1)~\cdots~\mathbf{x}(L)\big]$ in [rubin, liu2] and instead introduce the new complete data $(\mathbf{X},\mathbf{V})$ with $\mathbf{V}=\big[\mathbf{v}(1)~\cdots~\mathbf{v}(L)\big]$ [^3] The complete-data LLF of $(\mathbf{Y},\mathbf{X})$ is $l(\mathbf{X};\mathbf{S},\mathbf{Q})=\ln p(\mathbf{Y},\mathbf{X};\mathbf{S},\mathbf{Q})$ $={\sum}_{t=1}^{L}\Big[\ln p\big(\mathbf{y}(t)\mid\mathbf{x}(t);\mathbf{S},\mathbf{Q}\big)+\ln p\big(\mathbf{x}(t)\big)\Big]$ $=c-L\ln\mathrm{Det}(\mathbf{Q})-{\sum}_{t=1}^{L}\big[\mathbf{y}(t)-\mathbf{Sx}(t)\big]^{H}\mathbf{Q}^{-1}\big[\mathbf{y}(t)-\mathbf{Sx}(t)\big]$ where $c$ is a term unrelated to any free parameters in $(\mathbf{S},\mathbf{Q})$, and more complex than that of $(\mathbf{X},\mathbf{V})$ in (14).
+Numerical results show that the ECME algorithm with $(\mathbf{X},\mathbf{V})$ yields the same convergence as that with $(\mathbf{Y},\mathbf{X})$, but the ECME algorithm with $(\mathbf{X},\mathbf{V})$ needs fewer computations at each iteration..
+Toward this end, we rewrite the snapshots or measurements $\mathbf{Y}$ in (2) by [lawley, basilevsky, mulaik]
 $\mathbf{y}(t)=\mathbf{S}\mathbf{x}(t)+\mathbf{v}(t),t=1,\dots,L$ (13)
-where $\mathbf{x}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})$ is composed of $M$ common factors and uncorrelated with $\mathbf{v}(t)$. Then, the complete-data LLF is
-$l(\mathbf{X},\mathbf{V};\mathbf{Q})$ $=$ $\ln p(\mathbf{X})+\ln p(\mathbf{V};\mathbf{Q})$
+where $\mathbf{x}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})$ is composed of $M$ common factors and uncorrelated with $\mathbf{v}(t)$.
+Here, the statistical design of $\mathbf{x}(t)$ is based on $\mathbf{y}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{SS}^{H}+\mathbf{Q})$.
+Then, the complete-data LLF is
+$l(\mathbf{X},\mathbf{V};\mathbf{Q})$ $=$ $\ln p(\mathbf{X},\mathbf{V};\mathbf{Q})=\ln p(\mathbf{X})+\ln p(\mathbf{V};\mathbf{Q})$ (14)
 $=$ $c-L\Big(\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\{\hat{\mathbf{R}}_{v}\mathbf{Q}^{-1}\}\Big)$
-where $c$ is a term unrelated to any free parameters in $(\mathbf{S},\mathbf{Q})$, $\hat{\mathbf{R}}_{v}=(1/L)\sum_{t=1}^{L}\mathbf{v}(t)\mathbf{v}^{H}(t)$.
-With (4.1), we can construct the ECME algorithm, whose E-step and two CM-steps at the $k$th iteration are derived below.
+where $c$ is a term unrelated to any free parameters in $(\mathbf{S},\mathbf{Q})$ and $\hat{\mathbf{R}}_{v}=(1/L)\sum_{t=1}^{L}\mathbf{v}(t)\mathbf{v}^{H}(t)$.
+With (14), we can construct the ECME algorithm, whose E-step and two CM-steps at the $k$th iteration are derived below.
 4.1.1 First CM-step:
-Given $\mathbf{Q}=\mathbf{Q}^{(k-1)}$, let this step only obtain $\mathbf{S}^{(k)}$ by minimizing the actual or negative LLF $f(\mathbf{S},\mathbf{Q})$ with respect to $\mathbf{S}$, i.e.,
+Given $\mathbf{Q}=\mathbf{Q}^{(k-1)}\succ\mathbf{0}_{N}$, let this step only obtain $\mathbf{S}^{(k)}$ by directly minimizing the objective function $f(\mathbf{S},\mathbf{Q})$ in problem (6) with respect to $\mathbf{S}$, i.e.,
 $\min_{\mathbf{S}}\ln\mathrm{Det}\big(\mathbf{SS}^{H}+\mathbf{Q}^{(k-1)}\big)+\mathrm{Tr}\big\{\hat{\mathbf{R}}(\mathbf{SS}^{H}+\mathbf{Q}^{(k-1)})^{-1}\big\}.$ (15)
-This problem is classic [2, 4, 3] but the classic solutions in [18, 19] may not exist. Thus, we resort to the FAAN method by
+This problem is classic [lawley, basilevsky, mulaik] but the classic solutions in [joreskog, bartholomew] may not exist. Thus, we resort to the FAAN method by
 $\mathbf{S}^{(k)}=[\mathbf{Q}^{(k-1)}]^{1/2}\mathbf{U}^{(k)}[\boldsymbol{\Lambda}^{(k)}]^{1/2}\in\mathbb{C}^{N\times M},$ (16)
@@
 4.1.2 E-step:
-Given $\mathbf{Y}$, $\mathbf{S}^{(k)}$, and $\mathbf{Q}^{(k-1)}$, this step finds the expectation of $l(\mathbf{X},\mathbf{V};\mathbf{Q})$ in (4.1) by
-$l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ $=$ $\mathrm{E}\big{\{}l(\mathbf{X},\mathbf{V};\mathbf{Q})\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\}$
+Given $\mathbf{Y}$, $\mathbf{S}^{(k)}$, and $\mathbf{Q}^{(k-1)}$, this step finds the conditional expectation of the complete-data LLF $l(\mathbf{X},\mathbf{V};\mathbf{Q})$ in (14) by
+$l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})=\mathrm{E}\big\{l(\mathbf{X},\mathbf{V};\mathbf{Q})\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\}$ (17)
 $=c-L\Big(\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\{\hat{\mathbf{R}}_{v}^{(k)}\mathbf{Q}^{-1}\}\Big)$
-where $\boldsymbol{\Omega}^{(k)}=(\mathbf{S}^{(k)},\mathbf{Q}^{(k-1)})$, $\mathrm{E}\{\cdot\}$ and $\mathrm{D}\{\cdot\}$ means expectation and covariance, respectively. Moreover,
-$\hat{\mathbf{R}}^{(k)}_{v}=\mathrm{E}\big{\{}\hat{\mathbf{R}}_{v}\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}=[\boldsymbol{\delta}^{(k)}]^{H}\hat{\mathbf{R}}\boldsymbol{\delta}^{(k)}+\boldsymbol{\Delta}^{(k)}\succ\mathbf{0}_{N}$
-where $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}$, $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$ [^4] We do not adopt $\mathbf{C}^{(k)}]^{-1}=\mathbf{q}^{(k)}-\mathbf{q}^{(k)}\mathbf{S}^{(k)}(\boldsymbol{\Lambda}^{(k)}+\mathbf{I}_{M})^{-1}[\mathbf{S}^{(k)}]^{H}\mathbf{q}^{(k)}$ with $\mathbf{q}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1}$ since numerical results show its low accuracy., and $\boldsymbol{\Delta}^{(k)}=\mathbf{Q}^{(k-1)}-\mathbf{Q}^{(k-1)}\boldsymbol{\delta}^{(k)}\succeq\mathbf{0}_{N}$.
-The conditional distribution of $\mathbf{v}(t)$ can be derived from [20], and
-$\mathrm{E}\big{\{}\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}$ $=$ $[\boldsymbol{\delta}^{(k)}]^{H}\mathbf{y}(t),$
-$\mathrm{D}\big{\{}\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}$ $=$ $\boldsymbol{\Delta}^{(k)}\succeq\mathbf{0}_{N}.$
+where $\boldsymbol{\Omega}^{(k)}=(\mathbf{S}^{(k)},\mathbf{Q}^{(k-1)})$, $\mathrm{E}\{\cdot\}$ and $\mathrm{D}\{\cdot\}$ mean the expectation and covariance operators, respectively. Moreover,
+$\hat{\mathbf{R}}^{(k)}_{v}$ $=$ $\mathrm{E}\big\{\hat{\mathbf{R}}_{v}\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big\}$ (18)
+$=$ $\frac{1}{L}\sum_{t=1}^{L}\mathrm{E}\big\{\mathbf{v}(t)\mathbf{v}^{H}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big\}$
+$=$ $\frac{1}{L}\sum_{t=1}^{L}\Big(\boldsymbol{\Sigma}^{(k)}+\boldsymbol{\mu}^{(k)}(t)\big[\boldsymbol{\mu}^{(k)}(t)\big]^{H}\Big)$
+$=$ $\boldsymbol{\Sigma}^{(k)}+\frac{1}{L}\sum_{t=1}^{L}\boldsymbol{\mu}^{(k)}(t)[\boldsymbol{\mu}^{(k)}(t)]^{H}$
+$=$ $\boldsymbol{\Delta}^{(k)}+[\boldsymbol{\delta}^{(k)}]^{H}\hat{\mathbf{R}}\boldsymbol{\delta}^{(k)}\succ\mathbf{0}_{N}$
+where $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}\succ\mathbf{0}_{N}$, $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$ [^4] We do not adopt $\mathbf{C}^{(k)}]^{-1}=\mathbf{q}^{(k)}-\mathbf{q}^{(k)}\mathbf{S}^{(k)}(\boldsymbol{\Lambda}^{(k)}+\mathbf{I}_{M})^{-1}[\mathbf{S}^{(k)}]^{H}\mathbf{q}^{(k)}$ with $\mathbf{q}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1}$ since numerical results show its low accuracy., and $\boldsymbol{\Delta}^{(k)}=\mathbf{Q}^{(k-1)}-\mathbf{Q}^{(k-1)}\boldsymbol{\delta}^{(k)}\succeq\mathbf{0}_{N}$.
+Additionally, the conditional distribution of $\mathbf{v}(t)$ can be derived from [rhodes], and
+$\boldsymbol{\mu}^{(k)}(t)$ $=$ $\mathrm{E}\big\{\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big\}=[\boldsymbol{\delta}^{(k)}]^{H}\mathbf{y}(t),$ (19a)
+$\boldsymbol{\Sigma}^{(k)}$ $=$ $\mathrm{D}\big\{\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big\}=\boldsymbol{\Delta}^{(k)}\succeq\mathbf{0}_{N}.$ (19b)
 4.1.3 Second CM-step:
-This step only obtains $\mathbf{Q}^{(k)}$ by maximizing $l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ in (4.1.2) with respect to $\mathbf{Q}$, i.e.,
+This step only obtains $\mathbf{Q}^{(k)}$ by maximizing the expected complete-data LLF $l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ in (17) with respect to $\mathbf{Q}$, i.e.,
 $\min_{\mathbf{Q}\succ\mathbf{0}_{N}}\ln\mathrm{Det}(\mathbf{Q})+\mathrm{Tr}\big\{\hat{\mathbf{R}}_{v}^{(k)}\mathbf{Q}^{-1}\}.$ (20)
 This problem can be decomposed into the $N$ parallel subproblems:
-$\Big{(}\min_{\sigma_{n}^{2}>0}\ln(\sigma_{n}^{2})+[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}/\sigma^{2}_{n}\Big{)}~~n=1,\dots,N.$ (18)
-The solutions are
+$\min_{\sigma_{n}^{2}>0}\ln(\sigma_{n}^{2})+[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}/\sigma^{2}_{n},n=1,\dots,N.$ (21)
+Let $f(\sigma^{2}_{n})=\ln(\sigma^{2}_{n})+[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}/\sigma^{2}_{n}$ with $\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}>0$ and its derivative is
+$f^{\prime}(\sigma^{2}_{n})$ $=$ $\frac{1}{\sigma^{2}_{n}}-\frac{[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}}{\sigma^{4}_{n}}=\frac{\sigma^{2}_{n}-[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}}{\sigma^{4}_{n}}$ (25)
+[…]
+which indicates that when $\sigma^{2}_{n}=[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}$, $f(\sigma^{2}_{n})$ reaches the minimum. Thus, the solutions in subproblems (21) are
 $[\sigma^{2}_{n}]^{(k)}$ $=$ $[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}>0,n=1,\dots,N,$
 $\Rightarrow\mathbf{Q}^{(k)}$ $=$ $\mathrm{diag}\big(\hat{\mathbf{R}}_{v}^{(k)}\big)\succ\mathbf{0}_{N},$ (26)
-where $\mathrm{diag}(\cdot)$ is the matrix composed of the diagonal part. The details of the ECME algorithm are in Algorithm 2.
+where $\mathrm{diag}(\hat{\mathbf{R}}_{v}^{(k)})$ is the diagonal matrix composed of the diagonal part of $\hat{\mathbf{R}}_{v}^{(k)}$. The details of the ECME algorithm are in Algorithm 2.
 Algorithm 2 ECME Algorithm
@@
 4.2 Convergence and Complexity
-From [7, 21], we know that the ECME algorithm satisfies certain regularity conditions and its sequence $\big{\{}\big{(}\mathbf{S}^{(k)},\mathbf{Q}^{(k)}\big{)}\big{\}}$ always converges to a stationary point of $f(\mathbf{S},\mathbf{Q})$.
-Fortunately, simulation results in [10] and the subsequent section show that although the objective function $f(\mathbf{S},\mathbf{Q)}$ in problem (5) is non-convex, it seems to have a favourable landscape, which makes its minimization by the FAAN method or the ECME algorithm a relatively straightforward task.
-Moreover, Eq.
-(15) and Eq.
-(4.1.3) guarantee that the solution provided by the ECME algorithm is “proper" for problem (5) [18, 22].
-By comparing Algorithm 1 and Algorithm 2, we can find that at the $k$th iteration, the computational complexity of Algorithm 1 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (9) and the inverse of $\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}\in\mathbb{C}^{N\times N}$ in (3.2).
-Also, the computational complexity of Algorithm 2 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (15) and the inverse of $\mathbf{C}^{(k)}\in\mathbb{C}^{N\times N}$ in (4.1.3).
+From [liu, wu, McLachlan], we know that since only the complex Gaussian distribution, whose PDF has a regular exponential-family form, is involved, the ECME algorithm satisfies the “space-filling” and other regularity conditions and its sequence $\big\{\big(\mathbf{S}^{(k)},\mathbf{Q}^{(k)}\big)\big\}$ always converges to a stationary point of $f(\mathbf{S},\mathbf{Q})$.
+Theoretically, $f(\mathbf{S},\mathbf{Q})$ has multiple stationary points, so an accurate $\mathbf{Q}^{(0)}$ should be required by the ECME algorithm.
+Fortunately, simulation results in [stoica] and the subsequent section show that although the objective function $f(\mathbf{S},\mathbf{Q)}$ in problem (6) is non-convex, it seems to have a favourable landscape under moderately nonuniform noise, which makes its minimization by the FAAN method or the ECME algorithm a relatively straightforward task.
+However, since the second step at each iteration of the FAAN method sequentially (one by one) updates the estimates of $\sigma_{1},\dots,\sigma_{N}$ along the fixed coordinate directions, its sequence $\big\{\big(\mathbf{S}^{(k)},\mathbf{Q}^{(k)}\big)\big\}$ gets stuck at an undesired limit point more easily when the noise is immoderately nonuniform, which will be shown in Fig.
+3.
+We utilize the following proposition to guarantee that the solution provided by the ECME algorithm is “proper” for problem (6) [joreskog, adachi].
+Proposition 1.
+In the ECME algorithm, $\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}\succeq\mathbf{0}_{N}$ and $\mathbf{Q}^{(k)}\succ\mathbf{0}_{N}$ for arbitrary $k\geq 1$ if $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$ and $\mathbf{Q}^{(0)}\succ\mathbf{0}_{N}$.
+Proof.
+We utilize the mathematical induction method. When $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$ and $\mathbf{Q}^{(k-1)}\succ\mathbf{0}_{N}$, Eq. (16) directly indicates
+$\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}$
+$=[\mathbf{Q}^{(k-1)}]^{1/2}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}[\mathbf{Q}^{(k-1)}]^{1/2}\succeq\mathbf{0}_{N}$
+and then $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}\succ\mathbf{0}_{N}$ and $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$ are of full rank.
+As a result, $\boldsymbol{\delta}^{(k)}]^{H}\hat{\mathbf{R}}\boldsymbol{\delta}^{(k)}\succ\mathbf{0}_{N}\Rightarrow\hat{\mathbf{R}}^{(k)}_{v}=\boldsymbol{\Delta}^{(k)}+[\boldsymbol{\delta}^{(k)}]^{H}\hat{\mathbf{R}}\boldsymbol{\delta}^{(k)}\succ\mathbf{0}_{N}$ in (18) due to the covariance matrix $\boldsymbol{\Delta}^{(k)}\succeq\mathbf{0}_{N}$.
+Finally, we obtain $\mathbf{Q}^{(k)}=\mathrm{diag}\big(\hat{\mathbf{R}}^{(k)}_{v}\big)\succ\mathbf{0}_{N}$ in (4.1.3).
+The proof is completed.
+∎
+By comparing Algorithm 1 and Algorithm 2, Table 2 and Table 2 give their main operations of the $k$th iteration.
+From both tables, we can know that the computational complexities of each iteration in both algorithms involve $\mathit{\mathcal{O}}(N^{3})$ when $N$ is large [Izadkhah].
 Thus, the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method.
+Table 1: Computational complexity of the $k$th iteration in the FAAN method
+Main Operation eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}$ in (10) $\boldsymbol{\Gamma}^{(k)}=\big(\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big)^{-1}$ in (12) the $b_{n}^{(k)}$’s in (12)
+Complexity $\mathit{\mathcal{O}}(N^{3})$ $\mathit{\mathcal{O}}(N^{3})$ $\mathit{\mathcal{O}}(K_{2}N^{2})$
+Table 2: Computational complexity of the $k$th iteration in the ECME algorithm
+Main Operation eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}$ in (16) $\mathbf{C}^{(k)}]^{-1}$ in (18) matrix multiplications in (18)
+Complexity $\mathit{\mathcal{O}}(N^{3})$ $\mathit{\mathcal{O}}(N^{3})$ $\mathit{\mathcal{O}}(N^{3})$
 5 Numerical Results
 Simulation results are used to show and compare Algorithm 1 and Algorithm 2.
-Let $N=6$, $M=2$, $\mathbf{P}=10\times\mathbf{I}_{M}$, $\mathbf{Q}=\mathrm{diag}\{10,2,3,2,1,3\}$, $\theta_{1}=60^{\circ}$, $\theta_{2}=120^{\circ}$, and $K_{2}=100$.
-Moreover, $\mathbf{Q}^{(0)}=\mathbf{I}_{N}$.
+Unless otherwise specified in the following figures, let $N=6$, $M=2$, $\theta_{1}=60^{\circ}$, $\theta_{2}=120^{\circ}$, $\mathbf{P}=10\times\mathbf{I}_{M}$, $\mathbf{Q}=\mathrm{diag}\{10,2,3,2,1,3\}$, $K_{2}=100$, $L=100$, $K=K_{1}=100$, and $\mathbf{Q}^{(0)}=\mathbf{I}_{N}$.
+5.1 Convergence Evaluation
 Without loss of generality, Fig.
 1 illustrates the convergence performances of both algorithms given the same snapshots.
-Here, $L=100$.
+Here, the upper subfigure is with $\theta_{1}=90^{\circ}$ and $\theta_{2}=100^{\circ}$.
 We can clearly see that the ECME algorithm yields faster convergence than the FAAN method.
 Because the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method, it is computationally more efficient.
-Figure 1: Convergence comparison of the FAAN method and the ECME algorithm.
+Figure 1: Convergence comparison of the FAAN method and the ECME algorithm
 […]
+Fig.
+2 and Fig.
+3 show two scatter plots of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms under $100$ independent realizations.
+The same snapshots of each realization are processed by both algorithms.
+In Fig.
+2, the noise is moderately nonuniform and all the $(\hat{\theta}_{1},\hat{\theta}_{2})$’s are centered on the true value $(60^{\circ},120^{\circ})$ since for each realization, both sequences $\{\mathbf{S}^{(k)}\mathbf{S}^{(k)}\}$’s from the FAAN method and the ECME algorithm converge to the solution of $\mathbf{S}\mathbf{S}^{H}$ in problem (6).
+However, we know that under the same number of iterations $K=K_{1}$, both algorithms provide different $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$’s for each realization, and the $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$ provided by the ECME algorithm is closer to the solution of $\mathbf{S}\mathbf{S}^{H}$ in problem (6).
+In Fig.
+3, $\mathbf{Q}=\mathrm{diag}\{10,2,3000,2,1,3\}$, the numbers of desired points from the FAAN method and the ECME algorithm are $53$ and $98$, respectively.
+This implies that when the noise is immoderately nonuniform, the convergence of the FAAN method to an undesired limit point is easier.
+That is to say, the convergence of the ECME algorithm to the global optimum is easier.
+Figure 2: Scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms in moderately nonuniform noise
+Figure 3: Scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms in immoderately nonuniform noise
+When $\mathbf{P}=\begin{bmatrix}10~10\\ 10~10\end{bmatrix}$, i.e., the source signals are coherent, Fig.
+4 shows a scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms under $100$ independent realizations.
+The same snapshots of each realization are processed by both algorithms.
+In this figure, many $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms are not centered on the true value $(60^{\circ},120^{\circ})$ although the noise is moderately nonuniform.
+This indicates that in the presence of coherent source signals, the classic MLFA problem (6) cannot be employed directly to estimate the signal subspace and the spatial smoothing technique should be considered [Shan, Pillai].
+Figure 4: Scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms with coherent source signals
+5.2 Accuracy Evaluation
+In this subsection, the root mean square error (RMSE) performances of DOA estimation related to both algorithms are compared with the Cramer–Rao lower bound (CRLB) [pesavento, gershman].
+To guarantee the convergence of both algorithms to the global optimum, let the source signals be noncoherent and the noise be moderately nonuniform.
+In the following figures, each RMSE is based on $1000$ independent realizations and the RMSE of $\hat{\theta}_{m}$ is
+$\mathrm{RMSE}(\hat{\theta}_{m})$ $=$ $\sqrt{\frac{1}{1000}\sum_{i=1}^{1000}\Big(\hat{\theta}_{m}^{i}-\theta_{m}\Big)^{2}}$
+$=$ $10\lg\Big[\frac{1}{1000}\sum_{i=1}^{1000}\Big(\hat{\theta}_{m}^{i}-\theta_{m}\Big)^{2}\Big]~\mathrm{(dB)},$
+where $m=1,\dots,M$ and $\hat{\theta}_{m}^{i}$ denotes $\hat{\theta}_{m}$ obtained from the $i$th realization.
 […]
-Figure 2: Scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from the FAAN method and the ECME algorithm.
-Figure 3: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from the FAAN method and the ECME algorithm.
+Figure 5: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from both algorithms versus $L$
+Figure 6: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from both algorithms versus $\gamma$
+Figure 7: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2},\hat{\theta}_{3})$ from both algorithms versus $\gamma$
+Figure 8: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from both algorithms versus NN
+Fig.
+9 and Fig.
+10 compare the ECME algorithm with the state-of-the-art subspace separation approach in [Majdoddin], called “Good Method”.
+In Fig.
+9, $\mathbf{P}=\gamma\times\mathbf{I}_{M}$ while in Fig.
+10, $M=3$, $\theta_{1}=40^{\circ}$, $\theta_{2}=80^{\circ}$, $\theta_{3}=120^{\circ}$.
+We can observe that each RMSE curve of DOA estimation from the ECME algorithm is above that from the “Good Method”, so the noise subspace estimated by the “Good Method” is more accurate.
+Furthermore, because the “Good Method” only requires $5$ iterations and its computational complexity of each iteration also involves $\mathit{\mathcal{O}}(N^{3})$, it is computationally more efficient.
+As a consequence, MLFA is not the best choice for the subspace based DOA estimation in unknown nonuniform noise [^6] It is worth pointing out that the ECME algorithm is valuable since this algorithm should be the best choice for MLFA in other applications, e.g., psychometrics, econometrics, statistics and so on..
+Figure 9: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2})$ from the ECME algorithm and a method versus $\gamma$
+Figure 10: RMSE performances of $(\hat{\theta}_{1},\hat{\theta}_{2},\hat{\theta}_{3})$ from the ECME algorithm and a method versus $L$
 6 Conclusion
-In this article, the ECME algorithm, an extension of the EM algorithm, has been designed for MLFA, which has almost the same computational complexity at each iteration as the FAAN method with excellent convergence properties.
-However, numerical results have shown that the ECME algorithm yields faster stable convergence and is computationally more efficient.
-Importantly, the signal subspace estimated by the ECME algorithm can be employed for the subspace based DOA estimation in unknown nonuniform noise.
+In this article, the ECME algorithm, an extension of the EM algorithm, has been designed again for MLFA by introducing new complete data, which can thus use two explicit formulas to sequentially update the estimates of parameters at each iteration and have excellent convergence properties.
+Theoretical analysis has shown that the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method.
+However, numerical results have shown that the ECME algorithm yields faster stable convergence and the convergence to the global optimum is easier.
+Importantly, MLFA is not the best choice for the subspace based DOA estimation in unknown nonuniform noise.